### v1

# On Paszkiewicz’s conjecture about a product of positive contractions

Abstract.
The Paszkiewicz conjecture about a product of positive contractions asserts that given a decreasing sequence $T_{1}\geq T_{2}\geq\dots$ of positive contractions on a separable infinite-dimensional Hilbert space, the product $S_{n}=T_{n}\dots T_{1}$ converges strongly. Recently, the first named author verified the conjecture for certain classes of sequences. In this paper, we take a new approach by analysing the limit of the positive part $A=\lim_{n\to\infty}|S_{n}|$ (which always exists, regardless of the existence of $\lim_{n\to\infty}S_{n}$) and extend the class of examples for which the conjecture holds. We also show how an operator algebraic viewpoint can be useful to study the Paszkiewicz conjecture. Finally, we also show that the Paszkiewicz conjecture is true for all spectrally ordered sequences, i.e., those sequences for which $T_{1}^{k}\geq T_{2}^{k}\geq\dots$ holds for every $k\in\mathbb{N}$.
1. Introduction
Let $H$ be a separable infinite-dimensional Hilbert space, which we fix throughout the paper. The Paszkiewicz’s conjecture about a product of positive contraction is the following.
Conjecture 1.1 (Adam Paszkiewicz, 2018).
Let $H$ be a separable infinite-dimensional Hilbert space, $T_{1}\geq T_{2}\geq\dots$ be a sequence of positive linear contractions on $H$. Then the sequence $S_{n}:=T_{n}T_{n-1}\cdots T_{1}$ converges strongly.
Since $T_{1}\geq T_{2}\geq\dots$ is a decreasing sequence of positive contractions, the limit $T:=\lim_{n\to\infty}T_{n}$ (SOT) exists (SOT stands for the strong operator topology). We will use the notation that for a Borel subset $A$ of $\mathbb{R}$, $1_{A}(T)$ denotes the spectral projection of $T$ corresponding to $A$. Let $P:=1_{\{1\}}(T)$. In [2], Conjecture 1.1 is shown to be equivalent to Conjecture 1.2 below.
Conjecture 1.2.
Let $T_{1}\geq T_{2}\geq\dots$ be as in Conjecture 1.1. Then $\lim_{n\to\infty}S_{n}=P$ (*-strongly).
The Conjecture 1.2 is easily seen to be true in the following cases:
Example 1.3.
  • (1)
    The constant sequence $T_{n}\equiv T\,(\forall n)$. Then $S_{n}=T^{n}\to P$ (SOT).
  • (2)
    Each $T_{n}$ is a projection $P_{n}$. Then $T=\lim_{n}T_{n}$ is also a projection, say $P$, and $S_{n}=P_{n}\to P$ (SOT).
  • (3)
    $\|T_{n_{0}}\|<1$ for some $n_{0}$. Then because $\|T_{n+1}\|\leq\|T_{n}\|\leq\cdots$, $\|S_{n_{0}+k-1}\|\leq\|T_{n_{0}}\|^{k}\to 0\,(k\to\infty)$. Thus, $S_{n}\to P=0$ in norm.
  • (4)
    $T_{n}T_{m}=T_{m}T_{n}$ for all $n,m\in\mathbb{N}$. In this case, if $X$ is the Gelfand spectrum of the unital abelian C-algebra generated by $\{T_{1},T_{2},\dots\}$, then we may view $T_{n}=f_{n}$ for some $f_{n}\in C(X)$ and $f_{1}(x)\geq f_{2}(x)\geq\dots\geq f(x)=\lim_{n}f_{n}(x)$ for $x\in X$. Thus $S_{n}(x)=f_{n}(x)\cdots f_{1}(x)$. We may identify $H=L^{2}(X,\mu)$ for some Borel probability measure $\mu$ on $X$ with full support. Let $x\in X$. If $f(x)=1$, then $f_{k}(x)=1$ for every $k\in\mathbb{N}$ and thus $S_{n}(x)=1$ for every $n\in\mathbb{N}$. If $f(x)<1$, then there exists $k_{0}\in\mathbb{N}$ such that $f_{k}(x)<\frac{f(x)+1}{2}<1$ for every $k\geq k_{0}$, and thus $S_{n+k_{0}}(x)\leq\left(\frac{f(x)+1}{2}\right)^{n}\to 0\,(n\to\infty)$. Thus, $S_{n}(x)\to 1_{f^{-1}(\{1\})}$ pointwise, whence $S_{n}\to P$ (SOT).
Moreover, it is proved in [2] that $\lim_{n\to\infty}S_{n}^{*}=P$ (SOT), and the Paszkiewicz conjecture is true if either (i) the von Neumann algebra $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ generated by $T_{1},T_{2},\dots$ is finite, i.e., it admits a faithful normal tracial state, or (ii) $T_{1},T_{2},\dots$ has uniform spectral gap at 1, i.e., there exists $\delta\in(0,1)$ and $N\in\mathbb{N}$ such that $\sigma(T_{n})\cap(1-\delta,1)=\emptyset$ holds for all $n\geq N$. However, these are restrictive classes of sequences, and the general case remains out of reach.
In this paper, we extend the work [2], partly from operator algebraic viewpoint.
Below we explain the organization of the paper.
In §$\S$2, we focus on the positive part $|S_{n}|$ of $S_{n}$. It is easy to see that for any product $S_{n}=T_{n}\dots T_{1}$ of (not necessarily positive) contractions, its positive part $|S_{n}|$ converges in SOT. In the case of a decreasing sequence of positive contractions, we study the limit operator $A=\lim_{n\to\infty}|S_{n}|$, and in particular show in Theorem 2.7 that the Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if and only if $A$ is a projection, in which case $A=P$ holds. For each $k\in\mathbb{N}$, we have a positive contraction $A_{k}=\lim_{n\to\infty}|S_{n+k-1,k}|$, where $S_{n+k-1,k}=T_{n+k-1}\cdots T_{k}$. We show that the Paszkiewicz conjecture is true for $T_{1}\geq T_{2}\geq\dots$, if and only if $A_{k}=P$ for some $k$, if and only if $A_{k}=P$ for all $k$ (Proposition 2.10). Then we show that if $A_{1},A_{2},\dots$ has uniform spectral gap at 1, then the Paszkiewicz conjecture holds (Theorem 2.16). Note that the uniform spectral gap at 1 for $A_{1},A_{2},\dots$ is certainly necessary for the Paszkiewicz conjecture to be true, while the uniform spectral gap at 1 for $T_{1}\geq T_{2}\geq\dots$ need not hold for general sequences even when the Paszkiewicz conjecture holds for it.
In §$\S$3, we study the Paszkiewicz conjecture using operator algebraic viewpoint. From its formulation, it is not obvious that operator algebra theory can be useful to study the Paszkiewicz conjecture, but at least the conjecture can be verified if $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ is a finite von Neumann algebra by [2]. How large is the class of operators $T_{1},T_{2},\dots$ for which $\mathscr{M}$ is finite? Let us consider
Question 1.4.
How can one construct a decreasing sequence $T_{1}\geq T_{2}\geq\cdots$?
Is there some restriction on the structure of $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$?
If $x_{1},x_{2},\dots$ is a sequence of (not necessarily positive) contractions on $H$, then
$T_{n}=y_{n}^{*}y_{n},\,\,y_{n}=x_{n}\cdots x_{1}$
defines a decreasing sequence of positive contractions. Indeed, it is clear that for each $n\in\mathbb{N}$, $T_{n}$ is a positive contraction and
$T_{n+1}=x_{1}^{*}\cdots x_{n}^{*}x_{n+1}^{*}x_{n+1}x_{n}\cdots x_{1}\leq x_{1}^{*}\cdots x_{n}^{*}x_{n}\cdots x_{1}=T_{n}.$
We show that (i) any decreasing sequence of positive contractions arises this way (Proposition 3.1), and (ii) any von Neumann algebra $\mathscr{M}$ on $H$ is realized as $W^{*}(T_{1},T_{2},\dots)$ for some $T_{1}\geq T_{2}\geq\dots$ (Proposition 3.3). Thus, the finiteness assumption on $\mathscr{M}$ in [2] is indeed rather restrictive. Also, given that the only difficulty in proving the Paszkiewicz conjecture is the possibility of the SOT-discontinuity of the *-operation, the case where $\mathscr{M}$ is a type III factor seems interesting. On the other hand, we show that if $\mathscr{M}$ is a factor, then it is useful to consider an equivalent or a (possibly) stronger conjecture, which we explain below. The introduction of a more general conjecture at this stage might look strange, but it came from our experience that many of the arguments we have discovered so far to prove the original Paszkiewicz conjecture for some classes of sequences rely not too much on the fact that in the definition of the product $S_{n}=T_{n}\dots T_{1}$, the operators appear in the monotone decreasing order, though it is crucial that the $T_{n}$ converges to $T$. This leads us to consider the following reformulation/generalization of the Paszkiewicz conjecture. We denote by $\mathscr{S}$ (resp. $\mathscr{S}_{0}$) the set of all self-maps $\sigma\colon\mathbb{N}\to\mathbb{N}$ which are proper (resp. for which there exists $k\geq 0$ such that $\sigma(n+k)=n$ holds for all $n\in\mathbb{N}$). For $n\in\mathbb{N}$ and $\sigma\in\mathscr{S}$, define
$S_{n}^{\sigma}=T_{\sigma(n)}\dots T_{\sigma(1)}.$
Then one can show that the $\sigma$-Paszkiewicz subspace $H_{\sigma}=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}S_{n}^{\sigma}\xi=P\xi\right\}$ is a closed subspace of $H$, and we set
$H_{\mathscr{S}}=\bigcap_{\sigma\in\mathscr{S}}H_{\sigma},\,\,H_{\mathscr{S}_{0}}=\bigcap_{\sigma\in\mathscr{S}_{0}}H_{\sigma}.$
We say that the generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$, if $H_{\mathscr{S}}=H$ holds. Note that the Paszkiewicz conjecture for $T_{1}\geq T_{2}\geq\dots$ is equivalent to the condition $H_{\mathscr{S}_{0}}=H$. Obviously, the generalized Paszkiewicz conjecture implies the Paszkiewicz conjecture However, we show in Theorem 3.11 that if $\mathscr{M}$ is a factor, then the Paszkiewicz conjecture (resp. the generalized Paszkiewicz conjecture) holds if and only if $H_{\mathscr{S}_{0}}\neq\{0\}$ (resp. $H_{\mathscr{S}}\neq\{0\}$). The reason behind the above equivalence is the fact that the projection $e_{0}$ (resp. $e$) of $H$ onto $H_{\mathscr{S}_{0}}$ (resp. $H_{\mathscr{S}}$) is a central projection in $\mathscr{M}$. This provides an evidence that the operator algebraic viewpoint can be helpful for the study of the Paszkiewicz conjecture. We also remark that most of the examples of sequences we know so far for which the Paszkiewicz conjecture holds, do satisfy the generalized Paszkiewicz conjecture. In particular, we show that this is the case for sequences with uniform spectral gap at 1 (Theorem 3.12), extending the work [2].
Finally, we explain in §$\S$4 our earlier attempt to generalize the argument in [2] which worked to prove the Paszkiewicz conjecture for sequences with uniform spectral gap at 1. Though not entirely satisfactory, this attempt yielded as a by-product a new class of examples of sequences for which the Paszkiewicz conjecture is true. Namely, we show that if $T_{1}\succeq T_{2}\succeq\dots$ are spectrally ordered positive contractions, then the Paszkiewicz conjecture holds for it. Here for positive contractions $A,B$, we say $A$ is spectrally dominated by $B$ in the sense of Olson [7], denoted $A\preceq B$, if $1_{[c,1]}(A)\leq 1_{[c,1]}(B)$ for all $c\in(0,1)$, or equivalently if $A^{k}\leq B^{k}$ for all $k\in\mathbb{N}$ (see §$\S$4 for details). By Ogasawara’s theorem [6], a C-algebra $\mathcal{A}$ is commutative if and only if on the positive cone $\mathcal{A}_{+}$, the implication $0\leq A\leq B\implies 0\leq A^{2}\leq B^{2}$ holds. Note that since $t\mapsto t^{\alpha}$ is operator monotone if $\alpha\in(0,1]$, this implies $0\leq A^{k}\leq B^{k}$ for all $k\in\mathbb{N}$ and $A,B\in\mathcal{A}_{+}$ for which $A\leq B$ holds, i.e., $\leq$ and $\preceq$ agree on $\mathcal{A}_{+}$. However, it is known that the condition $A\preceq B$ alone for positive contractions $A,B$ does not necessarily imply that $AB=BA$. In fact, they can be highly non-commutative.
2. The limit operator $A=\lim_{n\to\infty}|S_{n}|$ and its truncated version $A_{1},A_{2},\dots$
We start from the following simple observation that for a product $S_{n}=T_{n}\cdots T_{1}$ of (not necessarily positive) contractions, its positive part $|S_{n}|$ always converges in SOT (this is most likely a folklore), even when the product itself fails to do so, e.g. as in the works of the Amemiya–Ando problem [4, 8, 5] on the SOT-convergence of random products of projections. On the other hand, there seems to be no reason to believe that $|S_{n}^{*}|$ converges for arbitrary product of contractions. However, in the case of a decreasing sequence of positive contractions, it converges to $P$.
Proposition 2.1.
Let $T_{1},T_{2},\dots$ be (not necessarily positive) contractions on $H$ and let $S_{n}=T_{n}\dots T_{1}$. Then its positive part $|S_{n}|$ converges in SOT to a positive contraction $A$. If moreover $T_{1}\geq T_{2}\geq\dots$ is a decreasing sequence of positive contractions on $H$. Then the following statements hold, with $P=1_{\{1\}}(T),\,T=\lim_{n\to\infty}T_{n}\,{\rm(SOT)}$.
  • (1)
    $\lim_{n\to\infty}|S_{n}^{*}|=P$ (SOT).
  • (2)
    $A=\lim_{n\to\infty}|S_{n}|$ (SOT) is a positive contraction satisfying $AP=P\leq A$.
Proof.
For each $n\in\mathbb{N}$, $0\leq T_{n+1}^{2}\leq 1$ implies
$S_{n+1}^{*}S_{n+1}$ $=T_{1}\dots T_{n}T_{n+1}T_{n+1}T_{n}\dots T_{1}$
$\leq T_{1}\dots T_{n}\cdot 1\cdot T_{n}\dots T_{1}=S_{n}^{*}S_{n}.$
This shows that $(S_{n}^{*}S_{n})_{n=1}^{\infty}$ is a decreasing sequence of positive contractions. Therefore $\lim_{n\to\infty}S_{n}^{*}S_{n}$ (SOT) exists, whence $A=\lim_{n\to\infty}|S_{n}|$ (SOT) exists.
Assume now that $T_{1}\geq T_{2}\geq\dots$ is a decreasing sequence of positive contractions.
(1) Let $\xi\in H$. Then by [2, Proposition 2.3 (1)], $\lim_{n\to\infty}S_{n}^{*}=P$ (SOT). It follows that
$\|(S_{n}S_{n}^{*}-P)\xi\|^{2}$ $=\|S_{n}S_{n}^{*}\xi\|^{2}+\|P\xi\|^{2}-2~{}{\rm Re}~{}\langle S_{n}S_{n}^{*}\xi,P\xi\rangle$
$\leq\|S_{n}^{*}\xi\|^{2}+\|P\xi\|^{2}-2~{}{\rm Re}~{}\langle S_{n}^{*}\xi,S_{n}^{*}P\xi\rangle$
$\xrightarrow{n\to\infty}2\|P\xi\|^{2}-2~{}{\rm Re}~{}\langle P\xi,P^{2}\xi\rangle=0.$
Therefore $\lim_{n\to\infty}S_{n}S_{n}^{*}=P$ (SOT), whence $\lim_{n\to\infty}|S_{n}^{*}|=P$ (SOT).
(2) Since $T_{n}P_{n}=P_{n}$ and $P\leq P_{n}$ ($P_{n}=1_{\{1\}}(T_{n})$) for every $n\in\mathbb{N}$ ([2, Lemma 2.1]), we have $T_{n}P=P$ for every $n\in\mathbb{N}$. This implies that $S_{n}^{*}S_{n}P=P$. Note that $|S_{n}|$ is the norm limit of a sequence of the form $(f_{k}(S_{n}^{*}S_{n}))_{k=1}^{\infty}$, where $f_{k}$ is a polynomial without constant term for each $k\in\mathbb{N}$. Therefore, it follows that $|S_{n}|P=P$ holds for every $n\in\mathbb{N}$. Thus, by letting $n\to\infty$ in the SOT, we obtain $AP=P$. Thus $P=AP=PA=APPA\leq A^{2}\leq A$ holds. ∎
In order to analyse the positive contraction $A$, let us introduce the following subspace of $H$, which we call the Paszkiewicz subspace. We remark that, inspired by the role played by the subspace “$Z$” in the work of Kopecká–Paszkiewicz [5, Lemma 3.4], we will introduce related closed subspaces $H_{\mathscr{S}_{0}}$ and $H_{\mathscr{S}}$ in §$\S$3. They will play a key role for an operator algebraic reformulation of the Paszkiewicz conjecture.
Proposition 2.2.
Let $T_{1}\geq T_{2}\geq\dots$ be positive contractions on $H$. Then the set
$\tilde{H}:=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}\|S_{n}\xi-P\xi\|=0\right\}=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}\|S_{n}\xi\|=\|P\xi\|\right\}$
is a closed subspace of $H$ containing $P(H)$.
Proof.
It is clear that $\tilde{H}$ is a vector subspace of $H$ and that $P(H)\subset\tilde{H}$ because $S_{n}P\xi=P\xi$ for every $\xi\in H$ and $n\in\mathbb{N}$. Let $\xi\in\overline{\tilde{H}}$ and $\varepsilon>0$. Then there exists $\xi_{0}\in\tilde{H}$ such that $\|\xi-\xi_{0}\|<\frac{\varepsilon}{3}$ holds. Since $\xi_{0}\in\tilde{H}$, we may find an $n_{0}\in\mathbb{N}$ such that $\|S_{m}\xi_{0}-S_{n}\xi_{0}\|<\frac{\varepsilon}{3}$ for every $n,m\geq n_{0}$. Then for every $n,m\geq n_{0}$,
$\|S_{m}\xi-S_{n}\xi\|$ $\leq\|S_{m}(\xi-\xi_{0})\|+\|S_{m}\xi_{0}-S_{n}\xi_{0}\|+\|S_{n}\xi_{0}-S_{n}\xi\|$
$<2\|\xi-\xi_{0}\|+\frac{\varepsilon}{3}<\varepsilon.$
Since $\varepsilon$ is arbitrary, this shows that $(S_{n}\xi)_{n=1}^{\infty}$ is Cauchy, whence it converges to $P\xi$ by [2, Proposition 2.3 (2)]. Therefore, $\xi\in\tilde{H}$ holds. This shows that $\tilde{H}$ is closed. Finally, to show the last equality, we show that for each $\xi\in H$ the following conditions are equivalent:
  • (a)
    $\lim_{n\to\infty}\|S_{n}\xi-P\xi\|=0.$
  • (b)
    $\lim_{n\to\infty}\|S_{n}\xi\|=\|P\xi\|.$
(a)$\implies$(b) is clear. Assume (b). Then because $\lim_{n\to\infty}S_{n}=P$ (WOT), we have
$\|S_{n}\xi-P\xi\|^{2}$ $=\|S_{n}\xi\|^{2}+\|P\xi\|^{2}-2\,{\rm Re}\langle S_{n}\xi,P\xi\rangle$
$\xrightarrow{n\to\infty}\|P\xi\|^{2}+\|P\xi\|^{2}-2\,{\rm Re}\langle P\xi,P\xi\rangle=0.$
Therefore (a) holds. This finishes the proof. ∎
Definition 2.3.
We call $\tilde{H}$ the Paszkiewicz subspace associated with $T_{1}\geq T_{2}\geq\dots$, and call the projection $Q$ onto $\tilde{H}$ the associated Paszkiewicz projection.
Proposition 2.4.
The positive contractions $A,Q,P\in\mathbb{B}(H)$ satisfy the following equalities.
  • (i)
    $\lim_{n\to\infty}S_{n}Q=P$ (SOT).
  • (ii)
    $1_{\{1\}}(A)=P$.
  • (iii)
    $AQ=AP=P$.
  • (iv)
    $1_{\{0\}}(A)\leq 1-P$.
Remark 2.5.
For (iv), note that proving the converse inequality
$1-P\leq 1_{\{0\}}(A)$
is equivalent to showing the Paszkiewicz conjecture for $T_{1}\geq T_{2}\geq\dots$.
In the proof Proposition 2.4 (ii), we will use the following lemma (see [1, Page 242, Step (ii) in the proof of Theorem]). The result holds more generally for contractions with the property (W) in the sense of [1].
Lemma 2.6 ([1]).
Let $T_{1},T_{2},\dots,T_{n}\in\mathbb{B}(H)$ be positive contractions, and let $S=T_{n}T_{n-1}\cdots T_{1}$. Let $\xi\in H$. Then
$\{\xi\in H\mid S\xi=\xi\}=\bigcap_{k=1}^{n}\{\xi\in H\mid T_{k}\xi=\xi\}.$
Proof of Proposition 2.4.
(i) is clear. (ii) By Proposition 2.1, $AP=P$ holds. In particular, $P\leq 1_{\{1\}}(A)$ holds. Conversely, if $\xi\in H$ satisfies $A\xi=\xi$, then $\lim_{n\to\infty}\|S_{n}\xi\|^{2}=\|\xi\|^{2}$ holds. Since $T_{1},T_{2},\dots$ are contractions, $\|\xi\|\geq\|S_{1}\xi\|\geq\|S_{2}\xi\|\geq\dots$ holds. Therefore, it holds that $\|S_{n}\xi\|=\|\xi\|$ for every $n\in\mathbb{N}$. Because the product of finitely many positive contractions satisfies the property (S) of [1, Page 240], hence a fortiori the property (W’) of [1], this implies that $S_{n}\xi=\xi$ for every $n\in\mathbb{N}$. Then it follows from Lemma 2.6, that $S_{n}\xi=\xi$ if and only if $T_{k}\xi=\xi$ for every $k=1,\dots,n$, if and only if $T_{n}\xi=1$ by $P_{1}\geq P_{2}\geq\dots\geq P_{n}$. Thus $T_{n}\xi=\xi$ for every $n\in\mathbb{N}$, whence $T\xi=\xi$, i.e., $\xi\in P(H)$ holds.
For the convenience of the reader, we also include the proof of $T_{n}\xi=\xi$, avoiding the introduction of (S) or (W’). First, $\|T_{1}\xi\|=\|\xi\|$ holds. By [2, Lemma 2.1], we have $T_{1}\xi=\xi$, and thus $\|S_{2}\xi\|=\|T_{2}\xi\|=\|\xi\|$, which by [2, Lemma 2.1] again implies $T_{2}\xi=\xi$. By induction, we have that $T_{n}\xi=\xi$ for every $n\in\mathbb{N}$.
(iii) by (ii), we may write $A=P+\tilde{A}$, where $\tilde{A}(=1_{[0,1)}(A))$ is a positive contraction satisfying $\tilde{A}P=P\tilde{A}=0$. Since $\lim_{n\to\infty}S_{n}^{*}=P$ (SOT), we have $\lim_{n\to\infty}QS_{n}^{*}=P$ (SOT). Then by (i), we have
$QA^{2}Q=\lim_{n\to\infty}QS_{n}^{*}S_{n}Q=P^{2}=P\,\,{\rm(SOT)}.$
On the other hand,
$QA^{2}Q=(P+(Q-P))(P+\tilde{A}^{2})(P+(Q-P))=P+(Q-P)\tilde{A}^{2}(Q-P).$
Therefore $(Q-P)\tilde{A}^{2}(Q-P)=0$, which implies that $\tilde{A}(Q-P)=0.$ Finally, we obtain $A(Q-P)=P(Q-P)+\tilde{A}(Q-P)=0$ by $PQ=P$. Therefore, $AQ=AP=P$ holds.
(iv) Let $\xi\in\ker(A)$. Then $\lim_{n\to\infty}\|S_{n}\xi\|=0$ holds. Let $n\in\mathbb{N}$. From the equality
$S_{n}\xi$ $=S_{n}P\xi+S_{n}P^{\perp}\xi$
$=P\xi+S_{n}P^{\perp}\xi,$
we see that since $S_{n}$ leaves $P^{\perp}(H)$ invariant, it follows that $P\xi=0$ and $\lim_{n\to\infty}\|S_{n}P^{\perp}\xi\|=0$ holds. Therefore, $\xi\in P^{\perp}(H)$ holds.
Now we show that $A$ is a projection precisely when the Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$.
Theorem 2.7.
Let $T_{1}\geq T_{2}\geq\dots$ be positive contractions on $H$, and let $A=\lim_{n\to\infty}|S_{n}|$ (SOT). The following three conditions are equivalent.
  1. (1)
    $\lim_{n\to\infty}S_{n}=P$ (SOT).
  2. (2)
    $A=P$.
  3. (3)
    $A$ is a projection.
Proof.
(2$\implies$(3) is clear, and (3$\implies$(2) follows from Proposition 2.4 (ii). Indeed, if $A$ is a projection, then we have $1_{\{1\}}(A)=P$ and $1_{\{0\}}(A)=1_{[0,1)}(A)=1_{\{1\}}(A)^{\perp}=P^{\perp}$.
(1$\implies$(2) It is clear that $A^{2}=\displaystyle\lim_{n\to\infty}S_{n}^{*}S_{n}=P^{2}=P$ because $\lim_{n\to\infty}S_{n}^{*}=P$ (SOT). Thus, we obtain $A=P$.
(2$\implies$(1) Let $S_{n}=u_{n}|S_{n}|$ be the polar decomposition of $S_{n}$. Since $S_{n}P=P$ and $|S_{n}|P=P$ for all $n\in\mathbb{N}$, we have $P=u_{n}P$. Hence,
$\|S_{n}\xi-P\xi\|=\|u_{n}|S_{n}|\xi-u_{n}P\xi\|$ $\leq\|u_{n}\|\||S_{n}|\xi-P\xi\|$
$\leq\||S_{n}|\xi-P\xi\|\xrightarrow{n\to\infty}0\quad(\xi\in H).$
So we obtain (1). ∎
2.1. Truncated product $S_{n,k}$
In view of Theorem 2.7, it is natural to expect $A=P$ for a large (potentially all) class of sequences $T_{1}\geq T_{2}\geq\dots$. The rest of this section is devoted to justifying this expectation in some cases. We introduce the following.
Definition 2.8.
Let $T_{1}\geq T_{2}\geq\dots$ be positive contractions on $H$. For each $n,k\in\mathbb{N}$ such that $n\geq k$, we define
$S_{n,k}=T_{n}T_{n-1}\cdots T_{k}$
Then by Proposition 2.1, the SOT-limit
$A_{k}=\lim_{n\to\infty}|S_{n,k}|$
exists. Since the sequence $T_{k}\geq T_{k+1}\geq\dots$ converges in SOT to $T$, we have by Proposition 2.1 that $\lim_{n\to\infty}|S_{n,k}^{*}|=P=1_{\{1\}}(T)$ (SOT).
The reason behind introducing these operators is the following. Since the Paszkiewicz conjecture is true for the constant case (Example 1.3), it is natural to expect that the conjecture could be true if the convergence of $T_{n}$ is fast, and because the Paszkiewicz conjecture is equivalent to $A=P$, it seems natural to expect that the limit of $S_{n,k}=T_{n}\cdots T_{k}$ as $n\to\infty$ is easier to establish when $k$ is fixed but large, and $A_{k}=\lim_{n\to\infty}|S_{n,k}|$ should be close to $P$ as $k\to\infty$. Thus, we state the following.
Question 2.9.
Let $T_{1}\geq T_{2}\geq\dots$ be a decreasing sequence of positive contractions on $H$. Is it true that $\lim_{k\to\infty}A_{k}=P$ (SOT)?
Though we have not been able to answer the above question completely, we have some partial results. First, the validity of the condition $A_{k}=P$ for one $k$ implies the validity of it for all $k$.
Proposition 2.10.
Let Let $T_{1}\geq T_{2}\geq\dots$ be a decreasing sequence of positive contractions on $H$. The following conditions are equivalent.
  • (i)
    $\lim_{n\to\infty}S_{n}=P$ (SOT).
  • (ii)
    $\lim_{n\to\infty}S_{n,k}=P$ (SOT) for some $k\in\mathbb{N}$.
  • (iii)
    $\lim_{n\to\infty}S_{n,k}=P$ (SOT) for every $k\in\mathbb{N}$.
Proof.
(iii)$\implies$(ii) is clear. Since $S_{n}=S_{n,k}T_{k-1}\cdots T_{1}$, (ii)$\implies$(i) holds. By an induction argument on $k$, for the proof of (i)$\implies$(iii), it suffices to show that $\lim_{n\to\infty}S_{n,2}=P$ (SOT). Let $\tilde{H}_{k}\,(k=1,2)$ be the Paszkiewicz subspace of $H$ associated with $T_{k}\geq T_{k+1}\geq\cdots$:
$\tilde{H}_{k}=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}S_{n,k}\xi=P\xi\right\},\,\,k=1,2.$
By (i), we have $\tilde{H}_{1}=H$. Since $S_{n,1}=S_{n,2}T_{1}$ and $\tilde{H}_{2}$ is closed by Proposition 2.2, it holds that $\overline{{\rm ran}}(T_{1})\subset\tilde{H}_{2}$ (if $\xi\in{\rm ran}(T_{1})$, then $\xi=T_{1}\eta$ for some $\eta\in H$. Therefore $S_{n,2}\xi=S_{n,1}\eta$, and because $\eta\in H=\tilde{H}_{1}$, $(S_{n}\eta)_{n=1}^{\infty}=(S_{n,2}\xi)_{n=1}^{\infty}$ converges, and the limit must be equal to $P\eta$ by [2, Proposition 2.3 (2)]. Then note that $P\xi=PT_{1}\eta=P\eta$). On the other hand, by $T_{1}\geq T_{2}$, we have $\ker{T_{1}}\subset\ker{T_{2}}\subset\cdots$. This implies that $\ker{T_{1}}\subset\ker{S_{n,2}}\subset\tilde{H}_{2}$. Therefore $H=\overline{{\rm ran}}(T_{1})\oplus\ker{T_{1}}\subset\tilde{H}_{2}$, whence $\lim_{n\to\infty}S_{n,2}=P$ (SOT) holds. ∎
2.2. The case of fast norm convergence of $T_{n}$
It seems likely to us that the answer to Question 2.9 is affirmative at least when $T_{n}$ converges to $T$ in norm. We verify $\lim_{k\to\infty}A_{k}=P$ under the norm convergence condition together with the additional hypothesis that if either (i) $\sum_{n=1}^{\infty}\|T_{n}-T\|<\infty$ or (ii) $1$ is isolated in $\sigma(T)$. We remark that actually, in these cases, we can give direct proofs of the Paszkiewicz conjecture (Propositions 2.13 and 2.15). Thus, the presentation here is somewhat redundant. We nevertheless include the arguments here, in a hope that arguments using the double limit theorem below could be useful for answering Question 2.9 in a more general setting.
Recall the following theorem (also called the Moore–Smith theorem) on the convergence of double limit:
Theorem 2.11 (Osgood’s double limit theorem).
Let $(X,d)$ be a complete metric space, $(x_{m,n})_{m,n=1}^{\infty}$ be a doubly indexed sequence in $X$. If
  • (i)
    $q_{m}=\lim_{n\to\infty}x_{m,n}$ exists for every $m\in\mathbb{N}$, and
  • (ii)
    $p_{n}=\lim_{m\to\infty}x_{m,n}$ exists for every $n\in\mathbb{N}$ and the convergence is uniform in $n$, i.e., $\lim_{m\to\infty}\sup_{n\in\mathbb{N}}d(x_{m,n},p_{n})=0$,
then $x:=\displaystyle\lim_{m\to\infty}q_{m}=\lim_{n\to\infty}p_{n}$ exists, and
$\lim_{m,n\to\infty}d(x_{m,n},x)=0.$
Proof.
Since we could not find a proper reference, we include the proof for completeness.
Step 1. We show that $(q_{m})_{m=1}^{\infty}$ is $d$-Cauchy.
Let $\varepsilon>0$. Then by (ii), there exists $m_{0}\in\mathbb{N}$ such that for every $n\in\mathbb{N}$ and every $m>m_{0}$, $d(x_{m,n},p_{n})<\frac{\varepsilon}{4}$ holds. Let $m,m^{\prime}\in\mathbb{N}$ be such that $m,m^{\prime}>m_{0}$. By (i), there exist $n_{m},n_{m^{\prime}}\in\mathbb{N}$ such that $d(x_{m,n},q_{m})<\frac{\varepsilon}{4}\,(n>n_{m})$ and $d(x_{m^{\prime},n},q_{m^{\prime}})<\frac{\varepsilon}{4}\,(n>n_{m^{\prime}})$. Let $n=\max(n_{m},n_{m^{\prime}})+1$. Then
$d(x_{m^{\prime},n},x_{m,n})$ $\leq d(x_{m^{\prime},n},p_{n})+d(p_{n},x_{m,n})$
$<\frac{\varepsilon}{2}.$
Therefore, we obtain
$d(q_{m},q_{m^{\prime}})$ $\leq d(q_{m},x_{m,n})+d(x_{m,n},x_{m^{\prime},n})+d(x_{m^{\prime},n},q_{m^{\prime}})$
$<\varepsilon.$
Since $\varepsilon$ is arbitrary, this shows that $(q_{m})_{m=1}^{\infty}$ is $d$-Cauchy. Thus by the completeness of $d$, the limit $x=\lim_{m\to\infty}q_{m}$ exists.
Step 2. We show that $\lim_{n\to\infty}p_{n}=x=\lim_{m,n\to\infty}x_{m,n}$ holds.
Let $\varepsilon>0$. Then by (ii), there exists $m_{1}\in\mathbb{N}$ such that $d(x_{m,n},p_{n})<\frac{\varepsilon}{6}$ holds for every $n\in\mathbb{N}$ and $m>m_{1}$. Fix $m>m_{1}$ for which $d(q_{m},x)<\frac{\varepsilon}{6}$ holds. By (i), there exists $n_{m}\in\mathbb{N}$ such that $d(x_{m,n},q_{m})<\frac{\varepsilon}{6}$ for every $n>n_{m}$. Then we have
$d(x_{m,n},x)$ $\leq d(x_{m,n},q_{m})+d(q_{m},x)<\frac{\varepsilon}{3}.$
Next, let $n,n^{\prime}\in\mathbb{N}$ be such that $n,n^{\prime}>n_{m}$. Then $d(x_{m,n},x_{m,n^{\prime}})\leq d(x_{m,n},x)+d(x,x_{m,n^{\prime}})<\frac{2\varepsilon}{3}$, whence (as before)
$d(p_{n},p_{n^{\prime}})$ $\leq d(p_{n},x_{m,n})+d(x_{m,n},x_{m,n^{\prime}})+d(x_{m,n^{\prime}},p_{n^{\prime}})$
$<\frac{\varepsilon}{6}+\frac{2\varepsilon}{3}+\frac{\varepsilon}{6}=\varepsilon.$
Since $\varepsilon>0$ is arbitrary, the limit $x^{\prime}=\lim_{n\to\infty}p_{n}$ exists. We show that $x=x^{\prime}$ holds. Let $\varepsilon>0$. Choose $m_{2}\in\mathbb{N}$ so that
(1) $d(x_{m,n},p_{n})<\varepsilon,\,\,\,n\in\mathbb{N},\,m>m_{2}$
Letting $n\to\infty$, we then obtain
$d(q_{m},x^{\prime})\leq\varepsilon\,(m>m_{1}).$
Then letting $m\to\infty$, we obtain
$d(x,x^{\prime})\leq\varepsilon.$
Since $\varepsilon>0$ is arbitrary, we see that $x=x^{\prime}$ holds. Finally, there exists $n_{1}$ such that $d(p_{n},x)<\varepsilon$ holds for every $n>n_{1}$. Then by (1), we have
$d(x_{m,n},x)<2\varepsilon,\,\,m>m_{2},\,n>n_{1}.$
Since $\varepsilon$ is arbitrary, we obtain $\lim_{m,n\to\infty}d(x_{m,n},x)=0$. ∎
Remark 2.12.
If $(x_{m,n})_{m,n=1}^{\infty}$ satisfies (i)(ii) above, then the convergence in (i) is uniform in $m$ as well. Indeed, let $\varepsilon>0$. Then there exist $m_{0},n_{0}\in\mathbb{N}$ such that $d(q_{m},x)<\frac{\varepsilon}{3}$ and $d(p_{n},x)<\frac{\varepsilon}{3}$ hold for every $m\geq m_{0}$ and $n\geq n_{0}$. Also by (ii), there exists $m_{1}\in\mathbb{N}$ such that $d(x_{m,n},p_{n})<\frac{\varepsilon}{3}$ for every $m\geq m_{1}$ and $n\in\mathbb{N}$. Moreover, by (i), there exists $n_{1}\in\mathbb{N}$ such that $d(x_{k,n},q_{k})<\varepsilon$ for every $k=1,\dots,m_{2}:=\max(m_{0},m_{1})$. Let $n_{2}=\max(n_{0},n_{1})$ and take arbitrary $n\geq n_{2}$ and $m\in\mathbb{N}$. If $m=1,\dots,m_{2}$, then $d(x_{m,n},q_{m})<\varepsilon$ holds. On the other hand, if $m>m_{2}$, then
$d(x_{m,n},q_{m})\leq d(x_{m,n},p_{n})+d(p_{n},x)+d(x,q_{m})<\varepsilon.$
Therefore, we have $d(x_{m,n},q_{m})<\varepsilon$ for every $m\in\mathbb{N}$. Since $\varepsilon$ is arbitrary, we see that the convergence in (i) is uniform in $m$.
We would like to apply Theorem 2.11 to $x_{m,n}=(S_{n+m-1,n})^{*}S_{n+m-1,n}$, where
$S_{n+m-1,n}=T_{n+m-1}\cdots T_{n},\,\,n,m\in\mathbb{N}.$
Note that
(2) $\lim_{m\to\infty}x_{m,n}=A_{n}^{2}\,\,{\rm(SOT)},$
and
(3) $\lim_{n\to\infty}x_{m,n}=T^{2m}\,\,{\rm(SOT).}$
Moreover, $\lim_{m\to\infty}T^{2m}=P$ (SOT). Note that the space $\mathbb{B}(H)_{1}$ of all contractions on $H$ is completely metrizable e.g., by a metric $d$ given by
$d(x,y)=\sum_{k=1}^{\infty}\frac{1}{2^{k}}\|(x-y)\xi_{k}\|,\,\,x,y\in\mathbb{B}(H)_{1},$
where $\{\xi_{k}\}_{k=1}^{\infty}$ is a dense subset of the unit ball of $H$. Then observe that the convergence in (2) is uniform in $n$ if and only if
(4) $\forall\xi\in H_{1}\,\forall\varepsilon>0\,\exists m_{0}\,\forall m\geq m_{0}\,\forall n\,\,\|(T_{n}\cdots T_{n+m-1}T_{n+m-1}\cdots T_{n}-A_{n}^{2})\xi\|<\varepsilon,$
and the convergence in (3) is uniform in $m$ if and only if
(5) $\forall\xi\in H_{1}\,\forall\varepsilon>0\,\exists n_{0}\,\forall n\geq n_{0}\,\forall m\,\,\|(T_{n}\cdots T_{n+m-1}T_{n+m-1}\cdots T_{n}-T^{2m})\xi\|<\varepsilon,$
Proposition 2.13.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$. Assume that $\sum_{k=1}^{\infty}\|T-T_{k}\|<\infty$ holds. Then $\lim_{k\to\infty}A_{k}^{2}=P$ (SOT).
Proof.
Let $n\in\mathbb{N}$. We have
$\|T_{n}\cdots$ $T_{n+m-1}T_{n+m-1}\cdots T_{n}-T^{2m}\|$
$\leq\sum_{k=n}^{n+m-2}\|T^{k-n}(T_{k}-T)T_{k+1}\cdots T_{n+m-1}T_{n+m-1}\cdots T_{n}\|$
$\hskip 28.45274pt+\|T^{m-1}(T_{n+m-1}-T)T_{n+m-1}\cdots T_{n}\|+$
$\hskip 28.45274pt+\sum_{k=n+1}^{n+m-1}\|T^{m}\cdot T^{n+m-k-1}(T_{k}-T)T_{k-1}\cdots T_{n}\|$
$\hskip 28.45274pt+\|T^{2m-1}(T_{n}-T)\|$
$\leq 2\sum_{k=n}^{n+m-1}\|T_{k}-T\|$
$\leq 2\sum_{k=n}^{\infty}\|T_{k}-T\|\xrightarrow{n\to\infty}0$
and the convergence is uniform in $m$. In particular, the convergence in (3) is uniform in $m$. Thanks to Theorem 2.11, it follows that
$\lim_{n\to\infty}A_{n}^{2}=\lim_{m\to\infty}T^{2m}=P\,\,{\rm(SOT)}.$
Actually, we have a direct proof of the Paszkiewicz conjecture when $\sum_{n=1}^{\infty}\|T-T_{n}\|<\infty$ holds.
Proposition 2.14.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$. Assume that $\sum_{n=1}^{\infty}\|T-T_{n}\|<\infty$ holds. Then the Paszkiewicz conjecture is true for $T_{1}\geq T_{2}\geq\dots$.
Proof.
Let $\xi\in H$ be a unit vector and $\varepsilon>0$. Choose $n_{0}\in\mathbb{N}$ such that $\sum_{n=n_{0}+1}^{\infty}\|T-T_{n}\|<\frac{\varepsilon}{2}$ holds. Since $\lim_{n\to\infty}T^{n}=P$ (SOT), there exists $n_{1}\in\mathbb{N}$ such that $\|T^{n}S_{n_{0}}\xi-PS_{n_{0}}\xi\|=\|T^{n}S_{n_{0}}\xi-P\xi\|<\frac{\varepsilon}{2}$ for every $n\geq n_{1}$.
Recall that $S_{n+n_{0},n_{0}+1}=T_{n+n_{0}}\cdots T_{n_{0}+1}$ ($n$ products). For every $n\in\mathbb{N}$, we have
$\|S_{n+n_{0},n_{0}+1}-T^{n}\|$ $\leq\sum_{k=1}^{n-1}\|T_{n+n_{0}}\cdots T_{n_{0}+k+1}(T_{n_{0}+k}-T)T^{k-1}\|+\|(T_{n+n_{0}}-T)T^{n-1}\|$
$\leq\sum_{k=n_{0}+1}^{\infty}\|T-T_{k}\|<\frac{\varepsilon}{2}.$
Therefore, by $S_{n+n_{0}}=S_{n+n_{0},n_{0}+1}S_{n_{0}}$, it holds that for every $n\geq n_{1}$,
$\|S_{n+n_{0}}\xi-P\xi\|$ $\leq\|(S_{n+n_{0},n_{0}+1}-T^{n})S_{n_{0}}\xi\|+\|T^{n}S_{n_{0}}\xi-PS_{n_{0}}\xi\|$
$<\|S_{n+n_{0},n_{0}+1}-T^{n}\|+\frac{\varepsilon}{2}<\varepsilon.$
Since $\varepsilon$ is arbitrary, this shows that $\lim_{n\to\infty}\|S_{n}\xi-P\xi\|=0$. ∎
Proposition 2.15.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$. Assume that $1$ is isolated in $\sigma(T)$ and $\lim_{n\to\infty}\|T_{n}-T\|=0$. Then $\lim_{n\to\infty}\|S_{n}-P\|=0$. In particular, the Paszkiewicz conjecture is true for $T_{1}\geq T_{2}\geq\cdots$.
Proof.
By assumption, there exists $\delta\in(0,1)$ such that $\sigma(T)\cap(1-\delta,1)=\emptyset$. In particular, $\|T^{k}-P\|=\|T^{k}P^{\perp}\|\leq(1-\delta)^{k}\xrightarrow{k\to\infty}0$. Therefore, for a given $\varepsilon>0$, there exists $k\in\mathbb{N}$ such that $\|T^{k}-P\|<\varepsilon$. Then by $T_{j}P=P\,(j\in\mathbb{N})$ and $\lim_{n\to\infty}\|T_{n+1}\dots T_{n+k}-T^{k}\|=0$, we have
$\limsup_{n\to\infty}\|S_{n}^{*}-P\|$ $=\limsup_{n\to\infty}\|S_{n+k}^{*}-P\|$
$=\limsup_{n\to\infty}\|T_{1}\dots T_{n}(T_{n+1}\dots T_{n+k}-P)\|$
$\leq\limsup_{n\to\infty}\|T_{n+1}\dots T_{n+k}-P\|$
$=\|T^{k}-P\|<\varepsilon.$
Since $\varepsilon$ is arbitrary, we obtain $\lim_{n\to\infty}\|S_{n}^{*}-P\|=\lim_{n\to\infty}\|S_{n}-P\|=0$. ∎
2.3. Uniform spectral gap at 1 for $A_{1},A_{2},\dots$ implies the Paszkiewicz conjecture
Recall that we have defined $A_{k}=\lim_{n\to\infty}|S_{n+k-1,k}|$ (SOT) and $A=\lim_{n\to\infty}|S_{n}|$ (SOT). The next theorem shows another evidence that the analysis of spectral properties of $A_{1},A_{2},\dots$ is relevant to the Paszkiewicz conjecture.
Theorem 2.16.
If $A_{1}=A,A_{2},\dots$ has uniform spectral gap at 1, then the Paszkiewicz conjecture holds for $T_{1},T_{2},\dots$.
Remark 2.17.
Note that if the Paszkiewicz conjecture is true for $T_{1},T_{2},\dots$ then $A_{k}=P$ for all $k$, whence $A_{1},A_{2},\dots$ clearly has uniform spectral gap at 1. This is in contrast to the hypothesis that $T_{1},T_{2},\dots$ has uniform spectral gap as in [2], which obviously need not be the case for general $T_{1},T_{2},\dots$.
The next lemma plays a key role in the proof of Theorem 2.16.
Lemma 2.18.
For every $\xi,\eta\in H$ and $k\in\mathbb{N}$, the following equality holds.
(6) $\langle A_{k}S_{k-1}\xi,A_{k}S_{k-1}\eta\rangle=\langle A\xi,A\eta\rangle.$
Here, we set $S_{0}:=1$. In particular, $\|A_{k}S_{k-1}\xi\|=\|A\xi\|$ holds.
Proof.
By definition, $A_{k}^{2}=\lim_{n\to\infty}S_{n+k-1,k}^{*}S_{n+k-1,k},A^{2}=\lim_{n\to\infty}S_{n}^{*}S_{n}$ (SOT) and $S_{n+k-1}=S_{n+k-1,k}S_{k-1}$. Therefore, we have
$\langle A_{k}S_{k-1}\xi,A_{k}S_{k-1}\eta\rangle$ $=\langle A_{k}^{2}S_{k-1}\xi,S_{k-1}\eta\rangle$
$=\lim_{n\to\infty}\langle S_{n+k-1,k}S_{k-1}\xi,S_{n+k-1,k}S_{k-1}\eta\rangle$
$=\lim_{n\to\infty}\langle S_{n}\xi,S_{n}\eta\rangle$
$=\langle A^{2}\xi,\eta\rangle=\langle A\xi,A\eta\rangle.$
Proof of Theorem 2.16.
Assume by contradiction that the Paszkiewicz conjecture fails for $T_{1},T_{2},\dots$. Then $A\neq P$ by Theorem 2.7, and therefore there exists $\xi\in P^{\perp}(H)$ for which $A\xi\neq 0$ holds. For each $n\in\mathbb{N}$, define $\tilde{T}_{n}=T_{n}P^{\perp}=P^{\perp}T_{n}$ and $\tilde{S}_{n}=S_{n}P^{\perp}=P^{\perp}S_{n}$. Then $\tilde{T}_{1}\geq\tilde{T}_{2}\geq\cdots$ is a decreasing sequence of positive contractions and $\tilde{S}_{n}=\tilde{T}_{n}\tilde{T}_{n-1}\cdots\tilde{T}_{1}=S_{n}P^{\perp}$. Thus $\tilde{A}_{k}=\lim_{n\to\infty}|\tilde{S}_{n+k-1,k}|=A_{k}P^{\perp}$ and $\tilde{A}=\lim_{n\to\infty}|\tilde{S}_{n}|=AP^{\perp}$. Since $A_{1},A_{2},\dots$ has uniform spectral gap at 1 and since $1_{\{1\}}(A_{k})=P\,(k\in\mathbb{N})$, there exist $N\in\mathbb{N}$ and $\delta\in(0,1)$ such that $\|\tilde{A}_{k}\|\leq\delta$ for all $k\geq N$.
On the other hand, we have
$\lim_{k\to\infty}\|\tilde{S}_{k-1}\xi\|=\|\tilde{A}\xi\|=\|A\xi\|\neq 0,$
and moreover by Lemma 2.18,
$\|A\xi\|=\|A_{k}S_{k-1}\xi\|=\|\tilde{A}_{k}\tilde{S}_{k-1}\xi\|,\,\,k\in\mathbb{N}.$
Then it follows that
11 11 $=\frac{\|\tilde{A}\xi\|}{\|\tilde{A}\xi\|}=\lim_{k\to\infty}\frac{\|\tilde{A}_{k}\tilde{S}_{k-1}\xi\|}{\|\tilde{S}_{k-1}\xi\|}$
$=\lim_{k\to\infty}\left\|\tilde{A}_{k}\left(\frac{\tilde{S}_{k-1}\xi}{\|\tilde{S}_{k-1}\xi\|}\right)\right\|$
$\leq\limsup_{k\to\infty}\|\tilde{A}_{k}\|\leq 1,$
which shows that $\limsup_{k\to\infty}\|\tilde{A}_{k}\|=1$, which contradicts $\|\tilde{A}_{k}\|\leq\delta\,(k\geq N)$. Therefore, the Paszkiewicz conjecture holds for $T_{1},T_{2},\dots$. ∎
3. Generalized Paszkiewicz conjecture: operator algebraic viewpoint
In this section, we take an operator algebra viewpoint to study the Paszkiewicz conjecture. We start from answering Question 1.4 in the introduction. If $x_{1},x_{2},\dots$ is a sequence of (not necessarily positive) contractions on $H$, then we have a sequence $T_{1}\geq T_{2}\geq\dots$ of positive contractions, where
$T_{n}=y_{n}^{*}y_{n},\,\,y_{n}=x_{n}\cdots x_{1},\,\,n\in\mathbb{N}.$
Actually, any $T_{1}\geq T_{2}\geq\dots$ is of this form:
Proposition 3.1.
Let $T_{1}\geq T_{2}\geq\cdots$ be a decreasing sequence of positive contractions on $H$. Then there exist a sequence $(x_{n})_{n=1}^{\infty}$ of contractions such that
$T_{n}=y_{n}^{*}y_{n},\,\,y_{n}=x_{n}\cdots x_{1},\,\,n\in\mathbb{N}.$
Moreover, $W^{*}(T_{1},T_{2},\dots)=W^{*}(x_{1},x_{2},\dots)$ holds.
The following lemma is well-known. We include the proof for completeness.
Lemma 3.2.
Let $S,T\in\mathbb{B}(H)$ be such that $0\leq S\leq T$. Then there exists a contraction $x\in W^{*}(S,T)$ such that $S^{\frac{1}{2}}=xT^{\frac{1}{2}}$.
Proof.
Set $\mathscr{M}=W^{*}(S,T)=W^{*}(S^{\frac{1}{2}},T^{\frac{1}{2}})$. Consider the orthogonal decomposition $H=\overline{\rm ran}(T^{\frac{1}{2}})\oplus\ker{T^{\frac{1}{2}}}$. We first define $x_{0}\colon{\rm ran}(T^{\frac{1}{2}})\to H$ by $x_{0}(T^{\frac{1}{2}}\xi)=S^{\frac{1}{2}}\xi$ for $\xi\in H$. Note that by $0\leq S\leq T$, we have $\|S^{\frac{1}{2}}\xi\|\leq\|T^{\frac{1}{2}}\xi\|$, so that $x_{0}$ is a well-defined contraction. Thus, it extends to a contraction from $\overline{\rm ran}(T^{\frac{1}{2}})$ to $H$, still denoted by $x_{0}$. We then define $x$ by setting $x_{0}$ on $\overline{\rm ran}(T^{\frac{1}{2}})$ and 0 on $\ker(T^{\frac{1}{2}})$. Then $S^{\frac{1}{2}}=xT^{\frac{1}{2}}$ holds. We show that $x\in\mathscr{M}$. Let $y^{\prime}$ be an element in the commutant ${M}^{\prime}$ of $\mathscr{M}$. For each $\xi\in H$, we have $y^{\prime}x_{0}(T^{\frac{1}{2}}\xi)=y^{\prime}S^{\frac{1}{2}}\xi=S^{\frac{1}{2}}y^{\prime}\xi=x_{0}(T^{\frac{1}{2}}y^{\prime}\xi)=x_{0}y^{\prime}(T^{\frac{1}{2}}\xi)$. Therefore, $y^{\prime}x_{0}=x_{0}y^{\prime}$ on $\overline{\rm ran}(T^{\frac{1}{2}})$. On the other hand, if $\xi\in\ker(T^{\frac{1}{2}})$, then $y^{\prime}\xi\in\ker(T^{\frac{1}{2}})$ by $T^{\frac{1}{2}}y^{\prime}=y^{\prime}T^{\frac{1}{2}}$, whence $y^{\prime}x\xi=xy^{\prime}\xi=0$. This shows that $y^{\prime}x=y^{\prime}x$. Therefore, we obtain $x\in\mathscr{M}^{\prime\prime}=\mathscr{M}$. ∎
Proof of Proposition 3.1.
Set $x_{1}=T_{1}^{\frac{1}{2}}$. For each $n\in\mathbb{N}$, we may apply Lemma 3.2 to $0\leq T_{n+1}\leq T_{n}$ to find a contraction $x_{n+1}\in W^{*}(T_{n+1},T_{n})$ such that $T_{n+1}^{\frac{1}{2}}=x_{n+1}T_{n}^{\frac{1}{2}}$. Then $T_{1}=x_{1}^{*}x_{1}$, and for $n\geq 2$,
$T_{n}$ $=T_{n-1}^{\frac{1}{2}}x_{n}^{*}x_{n}T_{n-1}^{\frac{1}{2}}$
$=T_{n-2}^{\frac{1}{2}}x_{n-1}^{*}x_{n}^{*}x_{n}x_{n-1}T_{n-2}^{\frac{1}{2}}$
$=\cdots=x_{1}^{*}\cdots x_{n}^{*}x_{n}\cdots x_{1}$
$=y_{n}^{*}y_{n},$
where $y_{n}=x_{n}\cdots x_{1}\,\,(n\in\mathbb{N})$. Finally, since $T_{n}=y_{n}^{*}y_{n}\in W^{*}(x_{1},x_{2},\dots)\,(n\in\mathbb{N})$, $W^{*}(T_{1},T_{2},\dots)\subset W^{*}(x_{1},x_{2},\dots)$ holds. On the other hand, $x_{1}=T_{1}^{\frac{1}{2}}\in W^{*}(T_{1},T_{2},\dots)$ and $x_{n+1}\in W^{*}(T_{n+1}^{\frac{1}{2}},T_{n}^{\frac{1}{2}})\subset W^{*}(T_{1},T_{2},\dots)\,(n\in\mathbb{N})$ implies that $W^{*}(x_{1},x_{2},\dots)\subset W^{*}(T_{1},T_{2},\dots)$. Therefore, $W^{*}(T_{1},T_{2},\dots)=W^{*}(x_{1},x_{2},\dots)$ holds. ∎
Next, we show that in general, there is no restriction on the possible structure of $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$.
Proposition 3.3.
Let $\mathscr{M}$ be a von Neumann algebra on $H$. Then there exist positive contractions $T_{1}\geq T_{2}\geq\dots$ such that $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ holds.
Proof.
Since $H$ is separable, there exists a countable family of projections $e_{1},e_{2},\dots$ in $\mathbb{B}(H)$ such that $\mathscr{M}=W^{*}(e_{1},e_{2},\dots)$. Then $x_{n}=\tfrac{1}{2}(e_{n}+1),\,(n\in\mathbb{N})$ is an invertible contraction and $W^{*}(x_{1},x_{2},\dots)=W^{*}(e_{1},e_{2},\dots)=\mathscr{M}$. Let $T_{n}=y_{n}^{*}y_{n}$, $y_{n}=x_{n}\cdots x_{1}$. Then $T_{1}\geq T_{2}\geq\cdots$ is a decreasing sequence of positive contractions on $H$ and ~$\tilde{\mathscr{M}}:=W^{*}(T_{1},T_{2},\dots)\subset\mathscr{M}$ holds. Since all $x_{1},x_{2},\cdots$ are invertible, so are $y_{1},y_{2},\dots$. We have $x_{1}=T_{1}^{\frac{1}{2}}\in\tilde{\mathscr{M}}$. Assume we have shown that $x_{1},\dots x_{n}\in\tilde{\mathscr{M}}$. Then $y_{1},\dots,y_{n}\in\tilde{\mathscr{M}}$, whence $T_{n+1}=y_{n}^{*}x_{n+1}^{2}y_{n}$ implies $x_{n+1}=((y_{n}^{*})^{-1}T_{n+1}y_{n}^{-1})^{\frac{1}{2}}\in\tilde{\mathscr{M}}$. This shows that $\mathscr{M}\subset\tilde{\mathscr{M}}$. Therefore, $\tilde{\mathscr{M}}=\mathscr{M}$ holds. ∎
We then consider an operator algebraic reformulation of the Paszkiewicz conjecture and its variant.
Definition 3.4.
A map $\sigma\colon\mathbb{N}\to\mathbb{N}$ is called proper, if for every $k\in\mathbb{N}$, the set $\sigma^{-1}(\{k\})$ is finite. The set of all proper maps from $\mathbb{N}$ to itself is denoted by $\mathscr{S}$. We say that $\sigma\in\mathscr{S}$ is almost an identity, if there exists $k\geq 0$ such that $\sigma(n+k)=n$ for every $n\in\mathbb{N}$. The set of all proper maps which are almost identity is denoted by $\mathscr{S}_{0}$.
Remark 3.5.
A map $\sigma\colon\mathbb{N}\to\mathbb{N}$ is proper if and only if $\lim_{n\to\infty}\sigma(n)=\infty$, i.e., for every $N\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that $\sigma(n)>N$ for every $n\geq n_{0}$.
Now for a proper map $\sigma\colon\mathbb{N}\to\mathbb{N}$, we set
$S_{n}^{\sigma}=T_{\sigma(n)}\dots T_{\sigma(1)}.$
Note that the truncated products $S_{n,k}$ (resp. $|A_{k}|$) in Definition 2.8 are of the form $S_{n-k+1}^{\sigma}$ (resp. $A_{\sigma}=\lim_{n\to\infty}|S_{n}^{\sigma}|$) for some $\sigma\in\mathscr{S}$.
By $\lim_{n\to\infty}\sigma(n)=\infty$, $\lim_{n\to\infty}T_{\sigma(n)}=T$ (SOT) holds. Then we have the following analogue of [2, Proposition 2.3]. The proof is essentially the same, so we do not repeat it here.
Proposition 3.6.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$, and let $S_{n}^{\sigma}:=T_{\sigma(n)}\cdots T_{\sigma(1)}$. The following statements hold (WOT stands for the weak operator topology):
  • (1)
    $\lim_{n\to\infty}(S_{n}^{\sigma})^{*}=P$ (SOT). In particular, $\lim_{n\to\infty}S_{n}^{\sigma}=P$ (WOT) holds.
  • (2)
    Let $\xi\in H$. If the set $\{S_{n}^{\sigma}\xi\mid\,n\in\mathbb{N}\}$ is totally bounded, then $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi-P\xi\|=0$ holds.
  • (3)
    For every $\xi\in H$ and every $k\in\mathbb{N}$, $\lim_{n\to\infty}\|S_{n+k}^{\sigma}\xi-S_{n}^{\sigma}\xi\|=0$ holds.
For each $\sigma\in\mathscr{S}$ and $n\in\mathbb{N}$, define $S_{n}^{\sigma}=T_{\sigma(n)}\cdots T_{\sigma(1)}$. Then $A_{\sigma}=\lim_{n\to\infty}|S_{n}^{\sigma}|$ (SOT) exists by Proposition 2.1, and also $\lim_{n\to\infty}(S_{n}^{\sigma})^{*}=\lim_{n\to\infty}|(S_{n}^{\sigma})^{*}|=P$ (SOT) by Proposition 3.6.
Definition 3.7.
We define the $\sigma$-Paszkiewicz subspace $H_{\sigma}$ by
$H_{\sigma}=\{\xi\in H\mid\lim_{n\to\infty}S_{n}^{\sigma}\xi=P\xi\},$
which is a closed subspace of $H$ (the same proof as in Proposition 2.2 using Proposition 3.6 (2) shows that $H_{\sigma}$ is indeed closed). We also define
$H_{\mathscr{S}}=\bigcap_{\sigma\in\mathscr{S}}H_{\sigma},\,\,H_{\mathscr{S}_{0}}=\bigcap_{\sigma\in\mathscr{S}_{0}}H_{\sigma}.$
Let $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$.
Remark 3.8.
Note that $P=1_{\{1\}}(T)\in\mathcal{Z}(\mathscr{M})$, because $T_{n}P=PT_{n}=P$ for every $n\in\mathbb{N}$, whence $P\in\mathscr{M}^{\prime}$, while $P=1_{\{1\}}(T)$ and $T=\lim_{n\to\infty}T_{n}\in\mathscr{M}$ implies $P\in\mathscr{M}$. Thus, if $\mathscr{M}$ is a factor and $P\neq 1$ (note that if $P=1$, then $T_{n}=1$ for all $n$ and the Paszkiewicz conjecture is trivially true for this sequence), then $P=0$.
Definition 3.9.
We say that
  • (1)
    The generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if $H_{\mathscr{S}}=H$ holds.
  • (2)
    The weak form of the Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if $\overline{\bigcup_{\sigma\in\mathscr{S}}H_{\sigma}}=H$ holds.
The next result is a corollary to Proposition 2.14.
Corollary 3.10.
Let $T_{1}\geq T_{2}\geq\dots$ be a norm-convergent sequence of positive contractions on $H$. Then there exists $\sigma\in\mathscr{S}$ such that $H_{\sigma}=H$. In particular, the weak form of the Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$.
Proof.
By $\lim_{n\to\infty}\|T_{n}-T\|=0$, there exist natural numbers $n_{1}<n_{2}<\dots$ such that $\|T_{n_{k}}-T\|<2^{-k}$ holds. Then $H_{\sigma}=H$ holds by Proposition 2.14, where $\sigma\in\mathscr{S}$ is defined by $\sigma(k)=n_{k},\,k\in\mathbb{N}$. ∎
Note that the Paszkiewicz conjecture for $T_{1}\geq T_{2}\geq\dots$ is equivalent to
(7) $H_{\mathscr{S}_{0}}=H.$
Theorem 3.11.
The following statements hold.
  • (1)
    Both $H_{\mathscr{S}}$ and $H_{\mathscr{S}_{0}}$ are closed subspaces of $H$ which are invariant under all $T_{n}$. In particular, the projection $e$ (resp. $e_{0}$) of $H$ onto $H_{\mathscr{S}}$ (resp. $H_{\mathscr{S}_{0}}$) belongs to $\mathscr{M}^{\prime}$.
  • (2)
    We have $e_{0}=z(e_{0})$ (resp. $e=z(e)$), where the right hand side is the central support of $e_{0}$ (resp. $e$) in $\mathscr{M}^{\prime}$. In particular, if $\mathscr{M}$ is a factor, then the Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if and only if $H_{\mathscr{S}_{0}}\neq\{0\}$.
Proof.
(1) Let $\sigma\in\mathscr{S}_{0}$ (resp. $\mathscr{S})$, $m,n\in\mathbb{N}$ and $\xi\in H_{\mathscr{S}_{0}}$. Then the map $\hat{\sigma}\colon\mathbb{N}\to\mathbb{N}$ defined by $\hat{\sigma}(1)=m,\hat{\sigma}(k+1)=\sigma(k),\,k\in\mathbb{N}$ is an element in $\mathscr{S}_{0}$ (resp. $\mathscr{S}$). Thus,
$S_{n}^{\sigma}(T_{m}\xi)=S_{n+1}^{\hat{\sigma}}\xi\xrightarrow{n\to\infty}P\xi=P(T_{m}\xi)$
by $\xi\in H_{\hat{\sigma}}$. Therefore $T_{m}\xi\in H_{\mathscr{S}_{0}}$ (resp. $H_{\mathscr{S}}$). This shows that $H_{\mathscr{S}_{0}}$ (resp. $H_{\mathscr{S}}$) is invariant under $T_{m}$. Therefore $e_{0}$ (resp. $e$) belongs to $\mathscr{M}^{\prime}$.
(2) Assume that $f_{0}$ is a projection in $\mathscr{M}^{\prime}$ such that $e_{0}\sim f_{0}$ in $\mathscr{M}^{\prime}$ ($\sim$ denotes the Murray-von Neumann equivalence of projections). By $e_{0}\sim f_{0}$, there exists a partial isometry $u\in\mathscr{M}^{\prime}$ such that $u^{*}u=e_{0}$ and $uu^{*}=f_{0}$. Let $\xi\in f_{0}(H)$ and $\sigma\in\mathscr{S}_{0}$. Then $u^{*}\xi\in e_{0}(H)\subset H_{\sigma}$. By $P,S_{n}^{\sigma}\in\mathscr{M}$ and $u\in\mathscr{M}^{\prime}$, we have
$S_{n}^{\sigma}\xi=S_{n}^{\sigma}uu^{*}\xi=uS_{n}^{\sigma}(u^{*}\xi)\xrightarrow{n\to\infty}uPu^{*}\xi=Puu^{*}\xi=P\xi.$
This shows that $\xi\in H_{\sigma}$. Since $\sigma\in\mathscr{S}_{0}$ is arbitrary, we obtain $\xi\in H_{\mathscr{S}_{0}}=e_{0}(H)$. Therefore, $f_{0}\leq e_{0}$ holds. It then follows that
$e_{0}\leq\bigvee\left\{ue_{0}u^{*}\mid u\in U(\mathscr{M}^{\prime})\right\}\leq\bigvee\left\{f_{0}\in{\rm Proj}(\mathscr{M}^{\prime})\mid f_{0}\sim e_{0}\right\}\leq e_{0},$
whence $e_{0}=\bigvee\left\{ue_{0}u^{*}\mid u\in U(\mathscr{M}^{\prime})\right\}=z(e_{0})$ holds. The proof of $e=z(e)$ is almost identical. The last claim is then immediate. ∎
We show that the generalized Paszkiewicz conjecture holds for most classes of sequences for which the Paszkiewicz conjecture has been verified. It is clear that the generalized Paszkiewicz conjecture is true when $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ is a finite von Neumann algebra, because $\lim_{n\to\infty}(S_{n}^{\sigma})^{*}=P$ (SOT) for every $\sigma\in\mathscr{S}$ and the *-operation is SOT-continuous on the unit ball of $\mathscr{M}$. The next result shows that the generalized Paszkiewicz conjecture is true also for sequences with uniform spectral gap at 1, which generalizes [2].
Proposition 3.12.
If $T_{1}\geq T_{2}\geq\dots$ has uniform spectral gap at 1, then the generalized Paszkiewicz conjecture holds for it.
Lemma 3.13.
Let $T_{1}\geq T_{2}$ be positive contractions on $H$. Let $P_{i}=1_{\{1\}}(T_{i})\,(i=1,2)$. Then for each $i,j\in\{1,2\}$, $T_{i}P_{j}^{\perp}(H)\subset P_{\max\{i,j\}}^{\perp}(H)$ holds.
Proof.
The case $i=j$ is trivial. Assume $i<j$. Then $T_{i}\geq T_{j}$, whence $P_{i}\geq P_{j}$ and $T_{i}P_{j}=P_{j}T_{i}=P_{j}$. Thus $T_{i}P_{j}^{\perp}(H)=P_{j}^{\perp}T_{i}(H)\subset P_{j}^{\perp}(H)$. If $i>j$, then $P_{j}\geq P_{i}$, whence $P_{j}^{\perp}\leq P_{i}^{\perp}$. Therefore, $T_{i}P_{j}^{\perp}(H)\subset T_{i}P_{i}^{\perp}(H)\subset P_{i}^{\perp}(H)$. ∎
Proof of Proposition 3.12.
Let $\sigma\in\mathscr{S}$. Let $\varepsilon>0$ and $\xi\in P^{\perp}(H)$ be a unit vector. Then there exists $n_{0}\in\mathbb{N}$ such that $\|\xi-P_{n_{0}}^{\perp}\xi\|<\frac{1}{2}\varepsilon$ holds. By assumption, there exists $N\in\mathbb{N}$ and $\delta\in(0,1)$ such that $\sigma(T_{n})\cap(1-\delta,1)=\emptyset$ for every $n\geq N$. We may assume that $N\geq n_{0}$. Using the properness of $\sigma$, we may find an increasing sequence of natural numbers
$2\leq n_{1}<n_{2}<\cdots$
such that $\sigma(n_{k})>\max\{N,\sigma(1),\dots,\sigma(n_{k}-1)\}$ for every $k\in\mathbb{N}$. Indeed, by the properness of $\sigma$, there exists $n_{1}\geq 2$ such that $\sigma(n_{1})\geq N$ holds. Assume that we have found $n_{1}<n_{2}<\cdots<n_{k}$. Then by the properness of $\sigma$, there exists $n\in\mathbb{N}$ for which $n>n_{k}$ and $\sigma(n)>\sigma(n_{k})$ holds. Let $n_{k+1}$ be the smallest such $n$. Then $\sigma(n_{k+1})>N,\sigma(n_{k})$, and if $n<n_{k}$, then $\sigma(n)<\sigma(n_{k})<\sigma(n_{k+1})$ and if $n_{k}\leq n<n_{k+1}$, then $\sigma(n_{k+1})>\sigma(n_{k})\geq\sigma(n)$ by the choice of $n_{k+1}$. By induction, we have the $n_{1}<n_{2}<\dots$ with the required properties. Choose $k\in\mathbb{N}$ for which $(1-\delta)^{k}<\frac{1}{2}\varepsilon$ holds. Then for every $n\geq n_{k}$, we have
$\|S_{n}^{\sigma}\xi\|\leq\|S_{n_{k}}^{\sigma}\xi\|\leq\|S_{n_{k}}^{\sigma}(\xi-P_{n_{0}}^{\perp}\xi)\|+\|S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
and by a repeated use of Lemma 3.13, we have
$S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi=T_{n_{k}}(\underbrace{T_{\sigma(n_{k}-1)}\cdots T_{\sigma(1)}P_{n_{0}}^{\perp}\xi)}_{\in P_{\max\{\sigma(1),\dots,\sigma(n_{k}-1),n_{0}\}}^{\perp}(H)}$
And by $\sigma(n_{k})>\max(\sigma(1),\dots,\sigma(n_{k}-1),n_{0},N)$, we have
$P_{\max\{\sigma(1),\dots,\sigma(n_{k}-1),n_{0}\}}^{\perp}(H)\subset P_{\sigma(n_{k})}^{\perp}(H)=1_{[0,1-\delta]}(T_{\sigma(n_{k})})(H).$
This implies that
$\|S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$ $\leq(1-\delta)\|T_{\sigma(n_{k}-1)}\cdots T_{\sigma(n_{k-1}+1)}S_{n_{k-1}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq(1-\delta)\|S_{n_{k-1}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq\dots\leq(1-\delta)^{k}\|S_{n_{1}-1}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq(1-\delta)^{k}<\frac{\varepsilon}{2}.$
Therefore $\|S_{n}^{\sigma}\xi\|<\varepsilon\,(n\geq n_{k})$. Since $\varepsilon$ is arbitrary, we obtain $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi\|=0$ for every $\xi\in P^{\perp}(H)$, i.e., $H_{\sigma}=H$ holds. ∎
Remark 3.14.
By Proposition 2.15, the generalized Paszkiewicz conjecture is also true if $\lim_{n\to\infty}\|T_{n}-T\|=0$ and $1\notin\sigma(T)$. Moreover, it is also easy to check that if $A_{1},A_{2},\dots$ are the operators that appear in the hypothesis of Theorem 2.16, then the operators $\tilde{A}_{k}$ in the proof of the theorem would satisfy $\limsup_{j\to\infty}\|\tilde{A}_{\sigma(k)}\|=1$ if we assume $H_{\sigma}\neq H$, whence we get by contradiction that $H_{\sigma}=H$. Thus, the generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ in this case as well.
4. Remark on the spectral order
In [2] the proof of the Paszkiewicz conjecture under the assumption of uniform spectral gap at 1 is based on the following two facts:
  • (i)
    $T_{1}\geq T_{2}\geq\cdots\searrow T$ implies $P_{1}\geq P_{2}\geq\dots\searrow P$ (SOT)
  • (ii)
    Each $T_{k}$ leaves the range of $P_{j}^{\perp}\,(k\geq j\gg 1)$ invariant.
Under the spectral gap condition, we have $P_{j}^{\perp}=1_{[1-\delta,1]}(T_{j})^{\perp}$ for $j\gg 1$. In the general case, one might hope to apply a similar technique by making an artificial spectral gap. Here is one such attempt we have made.
Proposition 4.1.
Let $T_{1}\geq T_{2}\geq\dots$ be a decreasing sequence of positive contractions on $H$ with $T=\lim_{n\to\infty}T_{n}{\rm(SOT)}$. Assume that there exists $\delta_{0}\in(0,1)$ such that for every $\delta\in(0,\delta_{0}]$, the following three conditions are satisfied. Then the Paszkiewicz conjecture is true for $T_{1}\geq T_{2}\geq\dots$.
  1. (1)
    $\lim_{n\to\infty}Q_{n,\delta}^{\perp}=Q_{\delta}^{\perp}$ (SOT), where $Q_{n,\delta}=1_{[1-\delta,1]}(T_{n})$ and $Q_{\delta}=1_{[1-\delta,1]}(T)$.
  2. (2)
    $Q_{n,\delta}^{\perp}\leq Q_{n+1,\delta}^{\perp}$ for every $n\in\mathbb{N}$.
  3. (3)
    For any $n\in\mathbb{N}$, $T_{m}$ leaves the range of $Q_{n,\delta}^{\perp}$ invariant whenever $m\leq n$.
For the proof, we need the next lemma. Its proof is straightforward, so we omit it.
Lemma 4.2.
Let $T_{1}\geq T_{2}\geq...\geq 0$ be a decreasing sequence of positive contractions on $H$ with $T=\lim_{n\to\infty}T_{n}{\rm(SOT)}$, and let $S_{n}=T_{n}\cdots T_{2}T_{1}$ for $n\in\mathbb{N}$. The following conditions are equivalent.
  • (1)
    The Paszkiewicz Conjecture holds for $T_{1}\geq T_{2}\geq\dots\geq 0$.
  • (2)
    There exists $\delta_{0}\in(0,1)$ such that for every $\delta\in(0,\delta_{0}]$, $\lim_{n\to\infty}S_{n}1_{[0,1-\delta)}(T)=0$ (SOT) holds.
Proof of Proposition 4.1.
By Lemma 4.2, it suffices to show that $S_{n}Q_{\delta}^{\perp}\to 0$ (SOT) as $n\to\infty$ for $\delta\in(0,\delta_{0}]$.
Let $\xi\in Q_{\delta}^{\perp}(H)$. Take $\varepsilon>0$. There exists $n_{0}\in\mathbb{N}$ s.t. $\|\xi-Q_{n,\delta}^{\perp}\xi\|<\varepsilon$ for all $n\geq n_{0}$ by (1). Also, we have $Q_{n_{0},\delta}^{\perp}\xi\in Q_{n_{0}+j,\delta}^{\perp}(H),j\in\mathbb{N}$ by (2). Thus, we obtain $S_{n_{0}+j-1}Q_{n_{0},\delta}^{\perp}\xi\in Q_{n_{0}+j,\delta}^{\perp}(H),j\in\mathbb{N}$ by (3).Therefore,
$\|S_{n_{0}+j}Q_{n_{0},\delta}^{\perp}\xi\|=\|T_{n_{0}+j}S_{n_{0}+j-1}Q_{n_{0},\delta}^{\perp}\xi\|$ $\leq\|T_{n_{0}+j}|_{Q_{n_{0}+j,\delta}^{\perp}(H)}\|\|S_{n_{0}+j-1}Q_{n_{0},\delta}^{\perp}\xi\|$
$\leq(1-\delta)\|T_{n_{0}+j-1}S_{n_{0}+j-2}Q_{n_{0},\delta}^{\perp}\xi\|$
$\leq...\leq(1-\delta)^{j}\|S_{n_{0}}Q_{n_{0},\delta}^{\perp}\xi\|,\quad j\in\mathbb{N}.$
Hence we have,
$\|S_{n_{0}+j}\xi\|$ $\leq\|S_{n_{0}+j}(\xi-Q_{n_{0},\delta}^{\perp}\xi)\|+\|S_{n_{0}+j}Q_{n_{0},\delta}^{\perp}\xi\|$
$\leq\|\xi-Q_{n,\delta}^{\perp}\xi\|+(1-\delta)^{j}\|S_{n_{0}}Q_{n_{0},\delta}^{\perp}\xi\|$
$<\varepsilon+(1-\delta)^{j}\|S_{n_{0}}Q_{n_{0},\delta}^{\perp}\xi\|,\quad j\in\mathbb{N}.$
Thus $\limsup_{j\to\infty}\|S_{n_{0}+j}\xi\|\leq\varepsilon$. Since $\varepsilon>0$ is arbitrary, we get $S_{n}\xi\to 0$ as $n\to\infty$. ∎
Let $T^{\prime}\leq T$ be positive contractions. By Proposition 4.1, we are led to consider when the implication $T^{\prime}\leq T\implies 1_{[1-\delta,1]}(T^{\prime})\leq 1_{[1-\delta,1]}(T)$ holds for sufficiently small $\delta$. We know that $1_{\{1\}}(T^{\prime})\leq 1_{\{1\}}(T)$ by [2, Corollary 2.2]. In other words, the Borel measurable function $1_{\{1\}}\colon[0,1]\to\mathbb{R}$ is operator monotone on $[0,1]$. However, it is well-known that for any $\delta\in(0,1)$, the Borel map $1_{[1-\delta,1]}\colon[0,1]\to[0,1]$ is not operator monotone. This follows from a general theorem due to Löwner. Indeed, by [3, Theorem 2.4.1], if $f\colon[0,1]\to\mathbb{R}$ is 2-monotone, then $f$ is C1 on $(0,1)$, but $1_{[1-\delta,1)}$ is not even continuous on $(0,1)$. Here, we include an explicit counter-example. This is extracted from the proof of [3, Theorem 2.4.1].
Example 4.3.
For $\delta\in(0,1)$,set
$A=\begin{bmatrix}\frac{1-\delta}{3}&0\\ 0&1-\delta\end{bmatrix},\,\,Q=\begin{bmatrix}\alpha_{1}&\sqrt{\alpha_{1}\alpha_{2}}\\ \sqrt{\alpha_{1}\alpha_{2}}&\alpha_{2}\end{bmatrix},\,\,B=A+Q,$
where $\alpha_{1}=\frac{4-\delta}{12},\alpha_{2}=\frac{\delta}{4}$. Then $I\geq B\geq A\geq 0$ but $1_{[1-\delta,1]}(B)\ngeq 1_{[1-\delta,1]}(A)$.
We define the divided difference
$f^{[1]}(a,b)=\frac{f(a)-f(b)}{a-b}\quad a,b\in\mathbb{R},\,a\neq b$
for a real valued function $f$.
Proposition 4.4.
Let $a,b\in\mathbb{R},\,a<b$. Assume that f is a real valued function on $(a,b)$, and let $\xi_{1},\xi_{2},\eta_{1},\eta_{2}\in\mathbb{R}$ be such that $a<\xi_{1}<\eta_{1}<\xi_{2}<\eta_{2}<b$. Set
$A=\begin{bmatrix}\xi_{1}&0\\ 0&\xi_{2}\end{bmatrix},\,Q=\begin{bmatrix}\alpha_{1}&\sqrt{\alpha_{1}\alpha_{2}}\\ \sqrt{\alpha_{1}\alpha_{2}}&\alpha_{2}\end{bmatrix},\,B=A+Q,$
where
$\alpha_{1}=\frac{(\eta_{1}-\xi_{1})(\eta_{2}-\xi_{1})}{\xi_{2}-\xi_{1}},\quad\alpha_{2}=\frac{(\xi_{2}-\eta_{1})(\eta_{2}-\xi_{2})}{\xi_{2}-\xi_{1}}.$
Then $B\geq A$ and eigenvalues of $B$ are $\eta_{1},\eta_{2}$. Also, $f(B)\geq f(A)$ implies
$\det\begin{bmatrix}f^{[1]}(\xi_{1},\eta_{1})&f^{[1]}(\xi_{1},\eta_{2})\\ f^{[1]}(\xi_{2},\eta_{1})&f^{[1]}(\xi_{2},\eta_{2})\end{bmatrix}\geq 0.$
Proof.
The proof is in [3, Theorem 2.4.1, step1 and step2]. ∎
Example 4.5.
Fix $10<\delta^{\prime}<\delta<$ 1.
Take $\xi_{1}\in(0,1-\delta),\eta_{1}\in(\xi_{1},1-\delta),\xi_{2}\in(1-\delta^{\prime},1),\eta_{2}\in(\xi_{2},1)$. Then $10<\xi_{1}<\eta_{1}<1-\delta<1-\delta^{\prime}<\xi_{2}<\eta_{2}<$ 1. Apply Proposition 4.4 for $\xi_{1},\xi_{2},\eta_{1},\eta_{2}$. We can easily calculate
$1^{[1]}_{[1-\delta,1]}(\xi_{1},\eta_{1})=1^{[1]}_{[1-\delta,1]}(\xi_{2},\eta_{2})=0,$
[…]
and then
$\det\begin{bmatrix}1^{[1]}_{[1-\delta,1]}(\xi_{1},\eta_{1})&1^{[1]}_{[1-\delta,1]}(\xi_{1},\eta_{2})\\ 1^{[1]}_{[1-\delta,1]}(\xi_{2},\eta_{1})&1^{[1]}_{[1-\delta,1]}(\xi_{2},\eta_{2})\end{bmatrix}<0.$
So we have $1_{[1-\delta,1]}(B)\ngeq 1_{[1-\delta,1]}(A)$.
On the other hand, we have $1_{[\varepsilon,1]}(A)=1_{[1-\delta,1]}(A)\;(\xi_{1}<\varepsilon\leq\xi_{2})$ since eigenvalues of $A$ are $\xi_{1},\xi_{2}$. In particular, $1_{[1-\delta^{\prime},1]}(A)=1_{[1-\delta,1]}(A)$ and hence $1_{[1-\delta,1]}(B)\ngeq 1_{[1-\delta^{\prime},1]}(A)$.
Therefore, for a given $T_{1}\geq T_{2}\geq\dots$, we cannot expect to have
(8) $1_{[1-\delta,1]}(T_{1})\geq 1_{[1-\delta,1]}(T_{2})\geq\dots$
for all $\delta\in(0,1)$. On the other hand, when we do have (8) for a specific $\delta$, and if $T$ is invertible, then $Q_{n}=1_{[1-\delta,1]}(T_{n})\to Q=1_{[1-\delta,1]}(T)$ (SOT) holds. We include its proof because it could still be useful even though $1_{[1-\delta,1]}$ is never operator monotone.
Proposition 4.6.
Let $T_{1}\geq T_{2}\geq...\geq 0$ be a decreasing sequence of positive contractions on $H$, and let $\delta\in(0,1)$.
Suppose that $0\notin\sigma(T)$, where $T=\lim_{n\to\infty}T_{n}$
(SOT). Then
$Q_{1,\delta}^{\perp}\leq...\leq Q_{n,\delta}^{\perp}\leq Q_{n+1,\delta}^{\perp}\leq...\leq Q_{\delta}^{\perp}\implies\displaystyle\lim_{n\to\infty}Q_{n,\delta}^{\perp}=Q_{\delta}^{\perp}{\rm(SOT)}.$
Lemma 4.7.
Let $T$ be a positive contraction on $H$, $\xi\in H$ and let $\delta\in(0,1)$. Then the following two conditions are equivalent.
  • (1)
    $\xi\in 1_{[1-\delta,1]}(T)H$.
  • (2)
    $\langle f_{k}(T)\xi,\xi\rangle\leq(1-\delta)^{-k}\|\xi\|^{2}$ for every $k\in\mathbb{N}$.
Here, $f_{k}\colon[0,1]\to\mathbb{R}$ is the (unbounded, discontinuous) Borel function given by
(9) $f_{k}(t)=\begin{cases}\,\,\,t^{-k}&(t\in(0,1])\\ \,\,\,\left(\frac{2}{1-\delta}\right)^{k}&(t=0)\end{cases}.$
(In particular, $f_{k}(T)$ is a possibly unbounded self-adjoint operator. Thus, $\langle f_{k}(T)\xi,\xi\rangle$ should be interpreted as $\infty$ if $\xi$ is not in the domain of $f_{k}(T)$)
Proof.
[…]
$\langle f_{k}(T)\xi,\xi\rangle$ $=\int_{[1-\delta,1]}t^{-k}d\langle E_{T}(t)\xi,\xi\rangle$
$\leq(1-\delta)^{-k}\int_{[1-\delta,1]}d\langle E_{T}(t)\xi,\xi\rangle$
$=(1-\delta)^{-k}\|\xi\|^{2}.$
(2$\implies$(1). Assume that $\xi\notin 1_{[1-\delta,1]}(T)H$. Then $1_{[0,1-\delta)}(T)\xi\neq 0$, and by $1_{[0,1-\delta)}(T)=\lim_{\varepsilon\to+0}1_{[0,1-\delta-\varepsilon]}(T)$ (SOT), there exists $\varepsilon\in(0,1-\delta)$ such that $\eta_{\varepsilon}=1_{[0,1-\delta-\varepsilon]}(T)\xi\neq 0$. Since $f_{k}(T)$ and $1_{[0,1-\delta-\varepsilon]}(T)$ are strongly commuting, it follows that
$\langle f_{k}(T)\xi,\xi\rangle$ $=\langle f_{k}(T)1_{[0,1-\delta-\varepsilon]}(T)\xi,\xi\rangle+\langle f_{k}(T)1_{(1-\delta-\varepsilon,1]}(T)\xi,\xi\rangle$
$\geq\left(\frac{2}{1-\delta}\right)^{k}\|1_{\{0\}}(T)\xi\|^{2}+\int_{(0,1-\delta-\varepsilon]}t^{-k}d\langle E_{T}(t)\eta_{\varepsilon}^{\prime},\eta_{\varepsilon}^{\prime}\rangle$
$\geq\left(\frac{2}{1-\delta}\right)^{k}\|1_{\{0\}}(T)\xi\|^{2}+(1-\delta-\varepsilon)^{-k}\|\eta_{\varepsilon}^{\prime}\|^{2}.$
Here, we set $\eta_{\varepsilon}^{\prime}=\eta_{\varepsilon}-1_{\{0\}}(T)\xi$. This implies that
(10) $\|1_{\{0\}}(T)\xi\|^{2}\leq\frac{1}{2^{k}}\|\xi\|^{2},\,\,k\in\mathbb{N},$
and
(11) $\|\eta_{\varepsilon}^{\prime}\|^{2}\leq\left(\frac{1-\delta-\varepsilon}{1-\delta}\right)^{k}\|\xi\|^{2},\,k\in\mathbb{N}.$
By (10), we have $01_{\{0\}}(T)\xi=$ 0, i.e., $\eta_{\varepsilon}^{\prime}=\eta_{\varepsilon}\neq 0$ holds. However, this is impossible because the left hand side of (11) is positive and is independent of $k$, while the right hand side tends to 0 as $k\to\infty$. ∎
Proof of Proposition 4.6.
There exist $\lim_{n\to\infty}Q_{n,\delta}=:Q^{\prime}_{\delta}\,{\rm(SOT)}$ because $Q_{1,\delta}\geq Q_{2,\delta}\geq...\geq Q_{\delta}\geq 0$. It is clear that $Q^{\prime}_{\delta}\geq Q_{\delta}$. We show that $Q^{\prime}_{\delta}\leq Q_{\delta}$.
Let $\xi\in Q^{\prime}_{\delta}(H)$. Then $\xi\in Q_{n,\delta}(H)$ for all $n\in\mathbb{N}$ because of the order of $\{Q_{n,\delta}\}_{n}$.
Let $f_{k}$ be the function defined in (9) of Lemma 4.7. Let $E_{T}(\cdot)$ be the spectral measure associated with $T$. Since $0\notin\sigma(T)$, there exists $\varepsilon>0$ such that $\sigma(T)={\rm{supp}}\,E_{T}\subseteq[\varepsilon,1]$. For $k\in\mathbb{N}$, set
$\tilde{f}_{k}(t)=\begin{cases}\,\,\,t^{-k}&(t\in[\varepsilon,1])\\ \,\,\,\varepsilon^{-k}&(t\in[0,\varepsilon))\end{cases}.$
We have $f_{k}(T)=\tilde{f}_{k}(T)$ because $f_{k}=\tilde{f}_{k}$ $E_{T}$-a.e. Since $\tilde{f}_{k}$ is continuous on $[0,1]$, we have $\tilde{f}_{k}(T_{n})\xrightarrow{n\to\infty}\tilde{f}_{k}(T)\,{\rm(SOT)}$. Then,
$\langle f_{k}(T)\xi,\xi\rangle=\langle\tilde{f}_{k}(T)\xi,\xi\rangle=\displaystyle\lim_{n\to\infty}\langle\tilde{f}_{k}(T_{n})\xi,\xi\rangle.$
Also, we have
$\langle\tilde{f}_{k}(T_{n})\xi,\xi\rangle=\int_{[1-\delta,1]}\tilde{f_{k}}(t)d\langle E_{T_{n}}(t)\xi,\xi\rangle\leq(1-\delta)^{-k}\|\xi\|^{2},\quad n\in\mathbb{N}$
because $\tilde{f_{k}}(1-\delta)\leq{\rm{min}}\{\varepsilon^{-k},(1-\delta)^{-k}\}$.
Therefore, we obtain $\langle f_{k}(T)\xi,\xi\rangle\leq(1-\delta)^{-k}\|\xi\|^{2}$. Since $k\in\mathbb{N}$ is arbitrary, we conclude by Lemma 4.7 that $\xi\in Q_{\delta}(H)$. Thus, $Q^{\prime}_{\delta}\leq Q_{\delta}$ holds, and then $Q_{\delta}=\displaystyle\lim_{n\to\infty}Q_{n,\delta}\,{\rm(SOT)}$ holds. ∎
Note that we do not need to have Eq. (8) for all $\delta\in(0,1)$ for Proposition 4.6. Only the condition (8) for sufficiently small $\delta$ is needed. We remark, however, that the condition
(12) $1_{[c,1]}(A)\leq 1_{[c,1]}(B)\text{\,\,\,for all\,\,\,}c\in(0,1)$
defines a partial ordering $\preceq$ among positive contractions. Such an ordering $\preceq$ is introduced by Olson [7] (for self-adjoint operators on a von Neumamnn algebra) and is called the spectral order. He showed the following equivalence about the spectral order.
Theorem 4.8 ([7, Theorem 3]).
Let $0\leq A\leq B$ be positive contractions. Then the following conditions are equivalent.
  • (i)
    $0\leq A^{k}\leq B^{k}$ for every $k\in\mathbb{N}$.
  • (ii)
    $1_{[1-\delta,1]}(A)\leq 1_{[1-\delta,1]}(B)$ for every $\delta\in(0,1)$.
We include the proof for the convenience of the reader.
Proof.
(i)$\implies$(ii) Let $\xi\in 1_{[1-\delta,1]}(A)$. In particular, $01_{\{0\}}(A)\xi=1_{\{0\}}(B)\xi=$ 0 holds. Let $f_{k}$ be as in Lemma 4.7. For each $m\in\mathbb{N}$, let $A_{m}=A+\frac{1}{m}$ and $B_{m}=B+\frac{1}{m}$. For each $k\in\mathbb{N}$ and $j=0,\dots,k$, we have $A^{j}\leq B^{j}$, so that
$A_{m}^{k}=\sum_{j=0}^{k}{k\choose j}m^{-(k-j)}A^{j}\leq\sum_{j=0}^{k}{k\choose j}m^{-(k-j)}B^{j}=B_{m}^{k}.$
Then we obtain $A_{m}^{-k}\geq B_{m}^{-k}$. Let $E_{A}$ (resp. $E_{B}$) be the spectral measure associated with $A$ (resp. $B$). Then $\langle A_{m}^{-k}\xi,\xi\rangle\geq\langle B_{m}^{-k}\xi,\xi\rangle$, i.e., we have
$\int_{(0,1]}\frac{1}{(t+\tfrac{1}{m})^{k}}d\langle E_{A}(t)\xi,\xi\rangle\geq\int_{(0,1]}\frac{1}{(t+\tfrac{1}{m})^{k}}d\langle E_{B}(t)\xi,\xi\rangle,$
which by the monotone convergence theorem and Lemma 4.7 for $A$ implies (use $01_{\{0\}}(A)\xi=1_{\{0\}}(B)\xi=$ 0)
$(1-\delta)^{-k}\|\xi\|^{2}\geq\langle f_{k}(A)\xi,\xi\rangle=\int_{(0,1]}t^{-k}d\langle E_{A}(t)\xi,\xi\rangle\geq\int_{(0,1]}t^{-k}d\langle E_{B}(t)\xi,\xi\rangle=\langle f_{k}(B)\xi,\xi\rangle.$
Therefore $\langle f_{k}(B)\xi,\xi\rangle\leq(1-\delta)^{-k}\|\xi\|^{2}$. Since $k\in\mathbb{N}$ is arbitrary, again by Lemma 4.7, we obtain $\xi\in 1_{[1-\delta,1]}(B)(H)$. This shows that $1_{[1-\delta,1]}(A)\leq 1_{[1-\delta,1]}(B)$.
(ii)$\implies$(i): For each $n\in\mathbb{N}$, define $h_{n}\colon[0,1]\to[0,1]$ by
$h_{n}(t)=\sum_{k=1}^{2^{n}-1}\frac{1}{2^{n}}1_{[\frac{k}{2^{n}},1]}(t),\,\,\,t\in[0,1].$
Then it holds that $0\leq x-h_{n}(x)\leq\frac{1}{2^{n}}$ for every $x\in[0,1]$. In particular, $h_{n}$ converges uniformly to the function $h\colon[0,1]\to[0,1]$ given by $h(x)=x\,(x\in[0,1])$. Therefore, we have $\lim_{n\to\infty}h_{n}(A)=A$ and $\lim_{n\to\infty}h_{n}(B)=B$ in SOT. Since $h_{n}(A)\leq h_{n}(B)$ for every $n\in\mathbb{N}$, it holds that $A\leq B$.
Let $k\in\mathbb{N}$ and $\delta\in(0,1)$. Then there is $\delta_{k}\in(0,1)$ for which $1-\delta_{k}=(1-\delta)^{\frac{1}{k}}$ holds, and then by (ii), we have $1_{[1-\delta,1]}(A^{k})=1_{[1-\delta_{k},1]}(A)\leq 1_{[1-\delta_{k},1]}(B)=1_{[1-\delta,1]}(B^{k})$. Thus, the condition (ii) with $A$ (resp. $B$) replaced by $A^{k}$ (resp. $B^{k}$) holds. Therefore, by what we have just shown, we obtain $A^{k}\leq B^{k}$. This shows (i). ∎
Finally, we show that the Paszkiewicz conjecture is true if the usual order is replaced by the spectral order.
Proposition 4.9.
Let $T_{1}\succeq T_{2}\succeq\dots$ be a decreasing sequence of spectrally ordered positive contractions on $H$. Then the Paszkiewicz conjecture holds for it.
Proof.
By Theorem 4.8, we have $T_{1}^{j}\geq T_{2}^{j}\geq\dots$ for every $j\in\mathbb{N}$. Let $k,n\in\mathbb{N}$. Then $\|S_{n+k}\xi\|=\|T_{n+k}T_{n+k-1}S_{n+k-2}\xi\|\leq\|T_{n+k-1}^{2}S_{n+k-2}\xi\|$ because $T_{n+k}^{2}\leq T_{n+k-1}^{2}$ implies $\|T_{n+k}\eta\|\leq\|T_{n+k-1}\eta\|$ for every $\eta\in H$. By an induction argument, we obtain
$\|S_{n+k}\xi\|$ $\leq\|T_{n+k-1}^{2}S_{n+k-2}\xi\|\leq\|T_{n+k-2}^{3}S_{n+k-3}\xi\|\leq\cdots\leq\|T_{k+1}^{n}S_{k}\xi\|.$
Letting $n\to\infty$, we obtain (use $T_{j}P_{k+1}=P_{k+1}=P_{k+1}T_{j}$ for $j=1,\dots,k$, where $P_{k+1}=1_{\{1\}}(T_{k+1})$)
$\|A\xi\|\leq\|P_{k+1}S_{k}\xi\|=\|P_{k+1}\xi\|.$
Then we let $k\to\infty$ to obtain
$\|A\xi\|\leq\|P\xi\|,$
which by $\|P\xi\|\leq\|A\xi\|$ implies $\|A\xi\|=\|P\xi\|$. Then by Proposition 2.2, $\xi$ belongs to the Paszkiewicz subspace for $T_{1},T_{2},\dots$. This finishes the proof. ∎
It is known that the spectral order $A\preceq B$ for positive contractions do not necessarily imply that $AB=BA$. Thus, the class of all spectrally ordered sequences is strictly larger than the class of sequences in which all operators commute. We include two examples to complete the picture.
Example 4.10.
Let $A=\begin{bmatrix}\frac{1}{3}&0\\ 0&0\end{bmatrix},B=\begin{bmatrix}\frac{5}{6}&\frac{1}{6}\\ \frac{1}{6}&\frac{5}{6}\end{bmatrix}$. Then it is easy to see that $A,B$ are positive contractions with spectral resolutions
$A=\tfrac{1}{3}\cdot\begin{bmatrix}1&0\\ 0&0\end{bmatrix}+0\cdot\begin{bmatrix}0&0\\ 0&0\end{bmatrix},\,\,\sigma(A)=\{0,\tfrac{1}{3}\},$
and
$B=\tfrac{2}{3}\cdot\begin{bmatrix}\tfrac{1}{2}&-\tfrac{1}{2}\\ -\tfrac{1}{2}&\tfrac{1}{2}\end{bmatrix}+1\cdot\begin{bmatrix}\tfrac{1}{2}&\tfrac{1}{2}\\ \tfrac{1}{2}&\tfrac{1}{2}\end{bmatrix},\,\,\sigma(B)=\{\tfrac{2}{3},1\},$
Moreover, $\sigma(B-A)=\{\tfrac{4-\sqrt{2}}{6},\tfrac{4+\sqrt{2}}{6}\}$ and $AB\neq BA$. We show that $A\preceq B$. Let $c\in(0,1)$.
If $c\in(0,\frac{1}{3}]$, then
$1_{[c,1]}(A)=\begin{bmatrix}1&0\\ 0&0\end{bmatrix}\leq 1=1_{[c,1]}(B).$
If $c\in(\tfrac{1}{3},\tfrac{2}{3}]$, then
$1_{[c,1]}(A)=0\leq 1=1_{[c,1]}(B).$
If $c\in(\tfrac{2}{3},1)$, then
$1_{[c,1]}(A)=0\leq\begin{bmatrix}\tfrac{1}{2}&\tfrac{1}{2}\\ \tfrac{1}{2}&\tfrac{1}{2}\end{bmatrix}=1_{[c,1]}(B).$
Therefore, $A\preceq B$ holds.
We also remark that the spectral order is closely related to commutativity as is clarified by the work of Uchiyama (see [9], in particular Theorem 1 therein). On the other hand, the next example shows that the condition $A\preceq B$ can be satisfied even when $A$ and $B$ are far from being commutative.
Example 4.11 (Communicated by Masaru Nagisa).
Let $A,C$ be positive operators such that $AC\neq CA$. Then $B=\|A\|1+C$ is a positive operator ($A,B$ can be made contractive by scaling) satisfying $A\preceq B$, but $AB\neq BA$. Indeed, for every $k\in\mathbb{N}$, we have
$B^{k}=(\|A\|1+C)^{k}\geq\|A\|^{k}+C^{k}\geq A^{k}.$
Acknowledgments
We would like to thank Professors Masaru Nagisa and Mitsuru Uchiyama for discussions regarding the spectral order and for pointing us to related literatures [7, 10] and Professor Nagisa for showing us Example 4.11. H. Ando is supported by Japan Society for the Promotion of Sciences KAKENHI 20K03647.