### v2

Abstract.
The Paszkiewicz conjecture about a product of positive contractions asserts that given a decreasing sequence $T_{1}\geq T_{2}\geq\dots$ of positive contractions on a separable infinite-dimensional Hilbert space, the product $S_{n}=T_{n}\dots T_{1}$ converges strongly. Recently, the first named author verified the conjecture for certain classes of sequences. In this paper, we prove the Paszkiewicz conjecture in full generality. Moreover, we show that in some cases, a generalized version of the Paszkiewicz conjecture also holds.
Mathematics Subject Classification (2020) 47A63, 47A65.
Keywords: positive contractions, operator products, Hilbert spaces, spectral theory
1. Introduction
Let $H$ be a separable infinite-dimensional Hilbert space, which we fix throughout the paper. The Paszkiewicz’s conjecture about a product of positive contraction is the following.
Conjecture 1.1 (Adam Paszkiewicz, 2018).
Let $T_{1}\geq T_{2}\geq\dots$ be a sequence of positive linear contractions on $H$. Then the sequence $S_{n}:=T_{n}T_{n-1}\cdots T_{1}$ converges strongly.
Since $T_{1}\geq T_{2}\geq\dots$ is a decreasing sequence of positive contractions, the limit $T:=\lim_{n\to\infty}T_{n}$ (SOT) exists (SOT stands for the strong operator topology). We will use the notation that for a Borel subset $A$ of $\mathbb{R}$, $1_{A}(T)$ denotes the spectral projection of $T$ corresponding to $A$. Let $P:=1_{\{1\}}(T)$. In [1], Conjecture 1.1 is shown to be equivalent to Conjecture 1.2 below.
Conjecture 1.2.
Let $T_{1}\geq T_{2}\geq\dots$ be as in Conjecture 1.1. Then $\lim_{n\to\infty}S_{n}=P$ (*-strongly).
The Conjecture 1.2 is easily seen to be true in the following cases:
Example 1.3.
  • (1)
    The constant sequence $T_{n}\equiv T\,(\forall n)$. Then $S_{n}=T^{n}\to P$ (SOT).
  • (2)
    Each $T_{n}$ is a projection $P_{n}$. Then $T=\lim_{n}T_{n}$ is also a projection, say $P$, and $S_{n}=P_{n}\to P$ (SOT).
  • (3)
    $\|T_{n_{0}}\|<1$ for some $n_{0}$. Then because $\|T_{n+1}\|\leq\|T_{n}\|\leq\cdots$, $\|S_{n_{0}+k-1}\|\leq\|T_{n_{0}}\|^{k}\to 0\,(k\to\infty)$. Thus, $S_{n}\to P=0$ in norm.
  • (4)
    $T_{n}T_{m}=T_{m}T_{n}$ for all $n,m\in\mathbb{N}$. In this case, if $X$ is the Gelfand spectrum of the unital abelian C-algebra generated by $\{T_{1},T_{2},\dots\}$, then we may view $T_{n}=f_{n}$ for some $f_{n}\in C(X)$ and $f_{1}(x)\geq f_{2}(x)\geq\dots\geq f(x)=\lim_{n}f_{n}(x)$ for $x\in X$. Thus $S_{n}(x)=f_{n}(x)\cdots f_{1}(x)$. We may identify $H=L^{2}(X,\mu)$ for some Borel probability measure $\mu$ on $X$ with full support. Let $x\in X$. If $f(x)=1$, then $f_{k}(x)=1$ for every $k\in\mathbb{N}$ and thus $S_{n}(x)=1$ for every $n\in\mathbb{N}$. If $f(x)<1$, then there exists $k_{0}\in\mathbb{N}$ such that $f_{k}(x)<\frac{f(x)+1}{2}<1$ for every $k\geq k_{0}$, and thus $S_{n+k_{0}}(x)\leq\left(\frac{f(x)+1}{2}\right)^{n}\to 0\,(n\to\infty)$. Thus, $S_{n}(x)\to 1_{f^{-1}(\{1\})}$ pointwise, whence $S_{n}\to P$ (SOT).
Moreover, it is proved in [1] that $\lim_{n\to\infty}S_{n}^{*}=P$ (SOT), and the Paszkiewicz conjecture is true if either (i) the von Neumann algebra $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ generated by $T_{1},T_{2},\dots$ is finite, i.e., it admits a faithful normal tracial state, or (ii) $T_{1},T_{2},\dots$ has uniform spectral gap at 1, i.e., there exists $\delta\in(0,1)$ and $N\in\mathbb{N}$ such that $\sigma(T_{n})\cap(1-\delta,1)=\emptyset$ holds for all $n\geq N$. However, these are restrictive classes of sequences. Indeed, any von Neumann algebra $\mathscr{M}$ on $H$ is of the form $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ for some decreasing sequence of positive contractions $T_{1}\geq T_{2}\geq\dots$ (Proposition 2.6). In this paper, we show that the Paszkiewicz conjecture is true in full generality.
In §$\S$2, we give a proof of the Paszkiewicz conjecture. We add an example to show that nevertheless, a product of decreasing sequence of positive contractions can have similar behaviour to orthogonal transformations (Proposition 2.1).
In §$\S$3, we consider a generalization of the Paszkiewicz conjecture. The motivation behind such a generalization came from our experience that many of our earlier arguments we have discovered so far to prove the original Paszkiewicz conjecture for some classes of sequences rely not too much on the fact that in the definition of the product $S_{n}=T_{n}\dots T_{1}$, the operators appear in the monotone decreasing order, though it is crucial that the $T_{n}$ converges to $T$. This leads us to consider the following generalization of the Paszkiewicz conjecture. We denote by $\mathscr{S}$ the set of all self-maps $\sigma\colon\mathbb{N}\to\mathbb{N}$ which are proper. For $n\in\mathbb{N}$ and $\sigma\in\mathscr{S}$, define
$S_{n}^{\sigma}=T_{\sigma(n)}\dots T_{\sigma(1)}.$
Then one can show that the $\sigma$-Paszkiewicz subspace $H_{\sigma}=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}S_{n}^{\sigma}\xi=P\xi\right\}$ is a closed subspace of $H$, and we set
$H_{\mathscr{S}}=\bigcap_{\sigma\in\mathscr{S}}H_{\sigma}.$
We say that the generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$, if $H_{\mathscr{S}}=H$ holds. We show in Theorem 3.7 that if $\mathscr{M}$ is a factor, then the generalized Paszkiewicz conjecture holds if and only if $H_{\mathscr{S}}\neq\{0\}$. We then remark that some classes of sequences do satisfy the generalized Paszkiewicz conjecture. In particular, we show that this is the case if either (1) $T_{1}$, $T_{2},\dots$ has uniform spectral gap at 1 (Theorem 3.8), extending the work [1], or (2) $\lim_{n\to\infty}\|T_{n}-T\|=0$ and $1$ is isolated in the spectrum $\sigma(T)$ of $T$ (the latter is a special case of the former, but it leads to the norm convergence of $S_{n}$. See Remark 3.11).
2. Proof of the Paszkiewicz conjecture
Here, we prove the Paszkiewicz conjecture (note that the seaparability of $H$ is not required).
Proof of the Paszkiewicz conjecture.
For each $m\in\mathbb{N}$, let $P_{m}=1_{\{1\}}(T_{m})$ and $P=1_{\{1\}}(T)$, with $T=\lim_{n\to\infty}T_{n}$ (SOT). Then $P_{1}\geq P_{2}\geq\dots\geq P$ and $\lim_{n\to\infty}P_{n}=P$ (SOT) by [1, Lemma 2.2]. Since $T_{n}P=PT_{n}=P$, we have $S_{n}P=P$ for each $n\in\mathbb{N}$. Thus, it suffices to show that $\lim_{n\to\infty}S_{n}P^{\perp}=0$ (SOT). By $P_{n}^{\perp}\nearrow P^{\perp}$ (SOT), it suffices to show that $S_{n}P_{m}^{\perp}\xrightarrow{n\to\infty}0$ (SOT) for every $m\in\mathbb{N}$.
Let $m\in\mathbb{N}$, $\xi_{1}\in P_{m}^{\perp}(H)$, and $\varepsilon>0$ be given and set $\xi_{n}:=T_{n-1}\xi_{n-1}=T_{n-1}\cdots T_{1}\xi_{1}$ for $n\geq 2$. It suffices to show $\lim_{n}\|\xi_{n}\|^{2}\leq\varepsilon$. Note that $\xi_{m}\in P_{m}^{\perp}(H)$ by $T_{j}P_{m}=P_{m}T_{j}=P_{m}\,(j\leq m)$, and that the sequence $\|\xi_{n}\|$ is decreasing. Consider the positive increasing functions $f_{k}(t):=1-(1-t)^{1/k}$ on $[0,1]$. Since $f_{k}$ converges to the characteristic function for $\{1\}$, one has
$\lim_{k\to\infty}\mathopen{\langle}f_{k}(T_{m})\xi_{m},\xi_{m}\mathclose{\rangle}=\|1_{\{1\}}(T_{m})\xi_{m}\|^{2}=0.$
Take $k$ such that $\mathopen{\langle}f_{k}(T_{m})\xi_{m},\xi_{m}\mathclose{\rangle}\leq\varepsilon/4$ and write $f:=f_{k}$. Since $f$ is operator monotone, one has
$\mathopen{\langle}f(T_{n})\xi_{n},\xi_{n}\mathclose{\rangle}\leq\mathopen{\langle}T_{n-1}f(T_{n-1})T_{n-1}\xi_{n-1},\xi_{n-1}\mathclose{\rangle}\leq\mathopen{\langle}f(T_{n-1})\xi_{n-1},\xi_{n-1}\mathclose{\rangle}\leq\cdots\leq\varepsilon/4$
[…]
$(1-\gamma^{2})\varepsilon/2\leq(1-\gamma^{2})\mu_{n}([0,\gamma))\leq\int 1-t^{2}\,d\mu_{n}(t)=\|\xi_{n}\|^{2}-\|\xi_{n+1}\|^{2}$
as long as $\|\xi_{n}\|^{2}\geq\varepsilon$. This implies $\|\xi_{n}\|^{2}\leq\varepsilon$ for any $n\geq m+2(1-\gamma^{2})^{-1}\varepsilon^{-1}\|\xi_{m}\|^{2}$. ∎
We complement the proof with the following example.
Proposition 2.1.
For every $10<\delta<$ 1, there are $n\in\mathbb{N}$ and a decreasing sequence $I\geq T_{1}\geq T_{2}\geq\cdots\geq T_{n}$ of positive contractions on the $2$-dimensional Hilbert space $\ell_{2}^{2}$ that satisfy $T_{n}\geq(1-\delta)I$ and
$T_{n}\cdots T_{1}\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)=\left(\begin{smallmatrix}0\\ 1-\delta\end{smallmatrix}\right)$
We need a preparatory construction. For a vector $\xi$, we denote by $P_{\xi}$ the rank-one orthogonal projection associated with $\xi$ and set $P_{\xi}^{\perp}:=I-P_{\xi}$. For a vector $\xi$, we denote by $P_{\xi}$ the rank-one orthogonal projection associated with $\xi$ and set $P_{\xi}^{\perp}:=I-P_{\xi}$.
Lemma 2.2.
Let $\varepsilon\in[0,1/2]$, $\theta\in[0,\pi/3]$, $\xi:=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, and $\eta:=\left(\begin{smallmatrix}\cos\theta\\ \sin\theta\end{smallmatrix}\right)$. Let $\phi\in[0,\pi/3]$, $\zeta=\left(\begin{smallmatrix}\cos\phi\\ \sin\phi\end{smallmatrix}\right)$, and $\alpha\in[0,1]$ be given by the equation
$(P_{\eta}+(1-\varepsilon)P_{\eta}^{\perp})\xi=\alpha\zeta.$
Then, one has $\alpha\geq 1-\varepsilon\theta^{2}$ and $|\phi-\varepsilon\theta|\leq C\varepsilon\theta^{3}$, where $C$ is an absolute constant.
Proof.
Write $c:=\cos\theta$ and $s:=\sin\theta\leq\theta$. Set $\eta^{\prime}:=\left(\begin{smallmatrix}s\\ -c\end{smallmatrix}\right)$. Then $\xi=c\eta+s\eta^{\prime}$ and
$(P_{\eta}+(1-\varepsilon)P_{\eta}^{\perp})\xi=c\eta+(1-\varepsilon)s\eta^{\prime}$
Hence
$\alpha=(1-2\varepsilon s^{2}+\varepsilon^{2}s^{2})^{1/2}$ $=((1-\varepsilon s^{2})^{2}+\varepsilon^{2}s^{2}(1-s^{2}))^{1/2}$
$=(1-\varepsilon s^{2})(1+t)^{1/2}\geq 1-\varepsilon\theta^{2},$
where $t:=\varepsilon^{2}s^{2}(1-s^{2})/(1-\varepsilon s^{2})^{2}=\varepsilon^{2}s^{2}+O(\varepsilon^{2}s^{4})$. This proves the first assertion. Since $(1+t)^{1/2}=1+t/2+O(t^{2})$ and $(1-u)^{-1}=1+u+O(u^{2})$,
$\alpha^{-1}=1+\varepsilon s^{2}-\frac{1}{2}\varepsilon^{2}s^{2}+O(\varepsilon^{2}s^{4}).$
Thus
$\cos\phi=\alpha^{-1}(1-\varepsilon s^{2})=1-\frac{1}{2}\varepsilon^{2}s^{2}+O(\varepsilon^{2}s^{4})=1-\frac{1}{2}\varepsilon^{2}\theta^{2}+O(\varepsilon^{2}\theta^{4}),$
by $s=\theta+O(\theta^{3})$. Since $0\leq\arccos(1-x)-(2x)^{1/2}\leq x^{3/2}$ for $x\in[0,1]$, one has
$\phi=\varepsilon\theta+O(\varepsilon\theta^{3}).$
This proves the second assertion (in fact for $C=10$ if the details are worked out). ∎
Lemma 2.3.
Let $20<\varepsilon<\varepsilon+\varepsilon^{2}<\kappa<1/$ 2, $\xi:=\left(\begin{smallmatrix}1\\ 0\end{smallmatrix}\right)$, and $\zeta:=\left(\begin{smallmatrix}\cos\phi\\ \sin\phi\end{smallmatrix}\right)$. Let $\beta\geq 0$ be the largest constant that satisfies
(1) $P_{\xi}+(1-\varepsilon)P_{\xi}^{\perp}\geq\beta(P_{\zeta}+(1-\kappa)P_{\zeta}^{\perp}).$
Then $1-2\phi^{2}\leq\beta\leq 1$.
Proof.
Note that (1) implies that $\beta P_{\zeta}$ is a positive contraction, whence $\beta\leq 1$ holds. We may exchange $\xi$ and $\zeta$.
Write $c:=\cos\phi$ and $s:=\sin\phi$. Thus, $\beta$ is the smaller of the eigenvalues of
$\left(\begin{smallmatrix}1&0\\ 0&(1-\kappa)^{-1/2}\end{smallmatrix}\right)\left(\begin{smallmatrix}1-\varepsilon s^{2}&\varepsilon cs\\ \varepsilon cs&1-\varepsilon c^{2}\end{smallmatrix}\right)\left(\begin{smallmatrix}1&0\\ 0&(1-\kappa)^{-1/2}\end{smallmatrix}\right)$ $=\left(\begin{smallmatrix}1-\varepsilon s^{2}&(1-\kappa)^{-1/2}\varepsilon cs\\ (1-\kappa)^{-1/2}\varepsilon cs&1+(1-\kappa)^{-1}(\kappa-\varepsilon+\varepsilon s^{2})\end{smallmatrix}\right).$
Since the eigenvalues of $\left(\begin{smallmatrix}a&b\\ b&d\end{smallmatrix}\right)$ are $((a+d)\pm((a-d)^{2}+4b^{2})^{1/2})/2$, we estimate for the above matrix
$a+d=2+\frac{\kappa-\varepsilon}{1-\kappa}+\frac{\kappa\varepsilon s^{2}}{1-\kappa}\geq 2+\frac{\kappa-\varepsilon}{1-\kappa}$
and, since $d-a=(1-\kappa)^{-1}(\kappa-\varepsilon+(2-\kappa)\varepsilon s^{2})\geq(1-\kappa)^{-1}(\kappa-\varepsilon)$,
$((a-d)^{2}+4b^{2})^{1/2}$ $\leq d-a+\frac{2b^{2}}{d-a}$
$\leq\frac{\kappa-\varepsilon}{1-\kappa}+\frac{(2-\kappa)\varepsilon s^{2}}{1-\kappa}+\frac{2\varepsilon^{2}c^{2}s^{2}}{\kappa-\varepsilon}$
$\leq\frac{\kappa-\varepsilon}{1-\kappa}+4s^{2}.$
Thus $\beta\geq 1-2s^{2}$. ∎
Proof of Proposition 2.1.
We may assume that $D:=2\delta^{-1}>8$ and set $\varepsilon_{k}:=(D+n-k)^{-1}$ and $\theta_{k}:=(\log(D+n-k))^{-1}\theta$ for $k=0,1,\ldots,n-1$, where $n\in\mathbb{N}$ and $\theta>0$ are chosen later. Note that $\varepsilon_{n}\leq\delta/2$. We recursively define $\xi_{k}:=\left(\begin{smallmatrix}\cos\phi_{k}\\ \sin\phi_{k}\end{smallmatrix}\right)$, $\eta_{k}:=\left(\begin{smallmatrix}\cos\rho_{k}\\ \sin\rho_{k}\end{smallmatrix}\right)$, and $\alpha_{k}$ as follows. First, set $\phi_{0}:=0$ and $\rho_{0}:=\theta_{0}$. For $k=0,\ldots,n-1$, set $\phi_{k+1}$ and $\alpha_{k+1}$ by the equation
$(P_{\eta_{k}}+(1-\varepsilon_{k})P_{\eta_{k}}^{\perp})\xi_{k}=\alpha_{k+1}\xi_{k+1}.$
Then, $\rho_{k+1}:=\phi_{k+1}+\theta_{k+1}$. Note that, by Lemma 2.2, one has
$\alpha_{k+1}\geq 1-\varepsilon_{k}\theta_{k}^{2}$
and
$|\phi_{k+1}-\phi_{k}-\varepsilon_{k}\theta_{k}|\leq C\varepsilon_{k}\theta_{k}^{3}.$
Note also that we may assume (by taking $\theta$ sufficiently small) that $C\varepsilon_{k}\theta_{k}^{3}\ll\varepsilon_{k}\theta_{k}$, and in particular that $C\varepsilon_{k}\theta_{k}^{3}\leq\frac{1}{2}\varepsilon_{k}\theta_{k}$ and $1-18\varepsilon_{k}\theta_{k}^{2}\geq\frac{1}{2}$ for all $k=0,\dots,n-1$. We recall that $\log(\log(t))^{\prime}=(t\log(t))^{-1}$. Thus,
$\phi_{n}\approx\sum_{k=0}^{n-1}\varepsilon_{k}\theta_{k}\approx\theta\log(\log n)$
as $n\to\infty$. Since $\phi_{n}$ depends continuously on $\theta$, by the intermediate value theorem, we may set $n$ and $\theta$ to be such that $\phi_{n}=\pi/2$, i.e., $\xi_{n}=\left(\begin{smallmatrix}0\\ 1\end{smallmatrix}\right)$. We use the inequality $t^{-1}(\log t)^{-2}\leq\frac{3}{2}(t+1)^{-1}(\log(t+1))^{-2},(t\geq 8)$ to obtain
00 0 $\leq\theta_{k+1}-\theta_{k}\leq(D+n-k-1)^{-1}(\log(D+n-k-1))^{-2}\theta$
$\leq\tfrac{3}{2}(D+n-k)^{-1}(\log(D+n-k))^{-2}\theta\leq\tfrac{3}{2}\varepsilon_{k}\theta_{k}.$
Thus, one has
00 0 $\leq\rho_{k+1}-\rho_{k}=\theta_{k+1}-\theta_{k}+\phi_{k+1}-\phi_{k}$
$\leq\tfrac{3}{2}\varepsilon_{k}\theta_{k}+\varepsilon_{k}\theta_{k}+C\varepsilon_{k}\theta_{k}^{3}$
$\leq 3\varepsilon_{k}\theta_{k}.$
Set $\beta_{k}>0$ to be the largest constant that satisfies
$(P_{\eta_{k-1}}+(1-\varepsilon_{k-1})P_{\eta_{k-1}}^{\perp})\geq\beta_{k}(P_{\eta_{k}}+(1-\varepsilon_{k})P_{\eta_{k}}^{\perp}).$
Since $\varepsilon_{k}-\varepsilon_{k-1}=(D+n-k)^{-1}(D+1+n-k)^{-1}>\varepsilon_{k-1}^{2}$, Lemma 2.3 implies
$\beta_{k}\geq 1-2(\rho_{k}-\rho_{k-1})^{2}\geq 1-18\varepsilon_{k-1}^{2}\theta_{k-1}^{2}.$
Finally, set $\beta_{0}:=1$ and
$T_{k}:=\beta_{0}\beta_{1}\cdots\beta_{k}(P_{\eta_{k}}+(1-\varepsilon_{k})P_{\eta_{k}}^{\perp})\leq T_{k-1}.$
Then, one has
$T_{n-1}\cdots T_{0}\xi_{0}=(\prod_{k=1}^{n}\alpha_{k})(\prod_{k=1}^{n-1}\beta_{k}^{n-k})\xi_{n}.$
For $B:=\int_{D}^{\infty}(t(\log t)^{2})^{-1}\,\mathrm{d}t=(\log D)^{-1}$,
$\sum_{k=1}^{n}(1-\alpha_{k})\leq\sum_{k=0}^{n-1}\varepsilon_{k}\theta_{k}^{2}\leq B\theta^{2}$
and by $\varepsilon_{k}(n-k)\leq 1$,
$\sum_{k=1}^{n-1}(1-\beta_{k})(n-k)\leq 18\sum_{k=0}^{n-2}\varepsilon_{k}^{2}\theta_{k}^{2}(n-k)\leq 18\sum_{k=0}^{n-1}\varepsilon_{k}\theta_{k}^{2}\leq 18B\theta^{2}.$
Since we arranged $\theta$ sufficiently small so that $0\leq 1-\alpha_{k}\leq\frac{1}{2}$ and $0\leq 1-\beta_{k}\leq\frac{1}{2}$ for all $k=1,\dots,n$ we have $\log(\alpha_{k})=\log(1-(1-\alpha_{k}))\geq-2(1-\alpha_{k})$ and $\log(\beta_{k})\geq-2(1-\beta_{k})$ thanks to the inequality $\log(1-x)\geq-2x\,(0\leq x\leq\frac{1}{2})$. Therefore,
$\sum_{k=1}^{n}\log\alpha_{k}$ $\geq\sum_{k=1}^{n}-2(1-\alpha_{k})\geq-2B\theta^{2},onehas$
and similarly
$\sum_{k=1}^{n-1}\log(\beta_{k}^{n-k})\geq-2\sum_{k=1}^{n-1}(n-k)(1-\beta_{k})\geq-32B\theta^{2}.$
These imply that $(\prod_{k=1}^{n}\alpha_{k})(\prod_{k=1}^{n-1}\beta_{k}^{n-k})\geq\exp(-34B\theta^{2})$.
Since $\theta>0$ could have been arbitrarily small, we may have $\exp(-34B\theta)>1-\delta$. Then replacing $T_{n}$ with $\lambda T_{n}$, where $\lambda\in(0,1)$ is chosen such that $\lambda(\prod_{k=1}^{n}\alpha_{k})(\prod_{k=1}^{n-1}\beta_{k}^{n-k})=1-\delta$, we are done. ∎
We end this section with the following remarks. First, if $x_{1},x_{2},\dots$ is a sequence of (not necessarily positive) contractions on $H$, then we have a decreasing sequence $T_{1}\geq T_{2}\geq\dots$ of positive contractions, where
$T_{n}=y_{n}^{*}y_{n},\,\,y_{n}=x_{n}\cdots x_{1},\,\,n\in\mathbb{N}.$
Actually, any $T_{1}\geq T_{2}\geq\dots$ is of this form:
Proposition 2.4.
Let $T_{1}\geq T_{2}\geq\cdots$ be a decreasing sequence of positive contractions on $H$. Then there exist a sequence $(x_{n})_{n=1}^{\infty}$ of contractions such that
$T_{n}=y_{n}^{*}y_{n},\,\,y_{n}=x_{n}\cdots x_{1},\,\,n\in\mathbb{N}.$
Moreover, $W^{*}(T_{1},T_{2},\dots)=W^{*}(x_{1},x_{2},\dots)$ holds.
The following lemma is well-known. We include the proof for completeness.
Lemma 2.5.
Let $S,T\in\mathbb{B}(H)$ be such that $0\leq S\leq T$. Then there exists a contraction $x\in W^{*}(S,T)$ such that $S^{\frac{1}{2}}=xT^{\frac{1}{2}}$.
Proof.
Set $\mathscr{M}=W^{*}(S,T)=W^{*}(S^{\frac{1}{2}},T^{\frac{1}{2}})$. Consider the orthogonal decomposition $H=\overline{\rm ran}(T^{\frac{1}{2}})\oplus\ker{T^{\frac{1}{2}}}$. We first define $x_{0}\colon{\rm ran}(T^{\frac{1}{2}})\to H$ by $x_{0}(T^{\frac{1}{2}}\xi)=S^{\frac{1}{2}}\xi$ for $\xi\in H$. Note that by $0\leq S\leq T$, we have $\|S^{\frac{1}{2}}\xi\|\leq\|T^{\frac{1}{2}}\xi\|$, so that $x_{0}$ is a well-defined contraction. Thus, it extends to a contraction from $\overline{\rm ran}(T^{\frac{1}{2}})$ to $H$, still denoted by $x_{0}$. We then define $x$ by setting $x_{0}$ on $\overline{\rm ran}(T^{\frac{1}{2}})$ and 0 on $\ker(T^{\frac{1}{2}})$. Then $S^{\frac{1}{2}}=xT^{\frac{1}{2}}$ holds. We show that $x\in\mathscr{M}$. Let $y^{\prime}$ be an element in the commutant $\mathscr{M}^{\prime}$ of $\mathscr{M}$. For each $\xi\in H$, we have $y^{\prime}x_{0}(T^{\frac{1}{2}}\xi)=y^{\prime}S^{\frac{1}{2}}\xi=S^{\frac{1}{2}}y^{\prime}\xi=x_{0}(T^{\frac{1}{2}}y^{\prime}\xi)=x_{0}y^{\prime}(T^{\frac{1}{2}}\xi)$. Therefore, $y^{\prime}x_{0}=x_{0}y^{\prime}$ on $\overline{\rm ran}(T^{\frac{1}{2}})$. On the other hand, if $\xi\in\ker(T^{\frac{1}{2}})$, then $y^{\prime}\xi\in\ker(T^{\frac{1}{2}})$ by $T^{\frac{1}{2}}y^{\prime}=y^{\prime}T^{\frac{1}{2}}$, whence $y^{\prime}x\xi=xy^{\prime}\xi=0$. This shows that $y^{\prime}x=y^{\prime}x$. Therefore, we obtain $x\in\mathscr{M}^{\prime\prime}=\mathscr{M}$. ∎
Proof of Proposition 2.4.
Set $x_{1}=T_{1}^{\frac{1}{2}}$. For each $n\in\mathbb{N}$, we may apply Lemma 2.5 to $0\leq T_{n+1}\leq T_{n}$ to find a contraction $x_{n+1}\in W^{*}(T_{n+1},T_{n})$ such that $T_{n+1}^{\frac{1}{2}}=x_{n+1}T_{n}^{\frac{1}{2}}$. Then $T_{1}=x_{1}^{*}x_{1}$, and for $n\geq 2$,
$T_{n}$ $=T_{n-1}^{\frac{1}{2}}x_{n}^{*}x_{n}T_{n-1}^{\frac{1}{2}}$
$=T_{n-2}^{\frac{1}{2}}x_{n-1}^{*}x_{n}^{*}x_{n}x_{n-1}T_{n-2}^{\frac{1}{2}}$
$=\cdots=x_{1}^{*}\cdots x_{n}^{*}x_{n}\cdots x_{1}$
$=y_{n}^{*}y_{n},$
where $y_{n}=x_{n}\cdots x_{1}\,\,(n\in\mathbb{N})$. Finally, since $T_{n}=y_{n}^{*}y_{n}\in W^{*}(x_{1},x_{2},\dots)\,(n\in\mathbb{N})$, $W^{*}(T_{1},T_{2},\dots)\subset W^{*}(x_{1},x_{2},\dots)$ holds. On the other hand, $x_{1}=T_{1}^{\frac{1}{2}}\in W^{*}(T_{1},T_{2},\dots)$ and $x_{n+1}\in W^{*}(T_{n+1}^{\frac{1}{2}},T_{n}^{\frac{1}{2}})\subset W^{*}(T_{1},T_{2},\dots)\,(n\in\mathbb{N})$ implies that $W^{*}(x_{1},x_{2},\dots)\subset W^{*}(T_{1},T_{2},\dots)$. Therefore, $W^{*}(T_{1},T_{2},\dots)=W^{*}(x_{1},x_{2},\dots)$ holds. ∎
Next, we show that any von Neumann algebra arises in the form of $W^{*}(T_{1},T_{2},\dots)$. Thus, the finiteness assumption of $W^{*}(T_{1},T_{2},\dots)$ in [1] is indeed quite restrictive.
Proposition 2.6.
Let $\mathscr{M}$ be a von Neumann algebra on $H$. Then there exist positive contractions $T_{1}\geq T_{2}\geq\dots$ such that $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ holds.
Proof.
Since $H$ is separable, there exists a countable family of projections $e_{1},e_{2},\dots$ in $\mathbb{B}(H)$ such that $\mathscr{M}=W^{*}(e_{1},e_{2},\dots)$. Then $x_{n}=\tfrac{1}{2}(e_{n}+1),\,(n\in\mathbb{N})$ is an invertible contraction and $W^{*}(x_{1},x_{2},\dots)=W^{*}(e_{1},e_{2},\dots)=\mathscr{M}$. Let $T_{n}=y_{n}^{*}y_{n}$, $y_{n}=x_{n}\cdots x_{1}$. Then $T_{1}\geq T_{2}\geq\cdots$ is a decreasing sequence of positive contractions on $H$ and ~$\tilde{\mathscr{M}}:=W^{*}(T_{1},T_{2},\dots)\subset\mathscr{M}$ holds. Since all $x_{1},x_{2},\cdots$ are invertible, so are $y_{1},y_{2},\dots$. We have $x_{1}=T_{1}^{\frac{1}{2}}\in\tilde{\mathscr{M}}$. Assume we have shown that $x_{1},\dots x_{n}\in\tilde{\mathscr{M}}$. Then $y_{1},\dots,y_{n}\in\tilde{\mathscr{M}}$, whence $T_{n+1}=y_{n}^{*}x_{n+1}^{2}y_{n}$ implies $x_{n+1}=((y_{n}^{*})^{-1}T_{n+1}y_{n}^{-1})^{\frac{1}{2}}\in\tilde{\mathscr{M}}$. This shows that $\mathscr{M}\subset\tilde{\mathscr{M}}$. Therefore, $\tilde{\mathscr{M}}=\mathscr{M}$ holds. ∎
3. Generalization of the Paszkiewicz conjecture
In this section, we consider a generalization of the Paszkiewicz conjecture.
Definition 3.1.
A map $\sigma\colon\mathbb{N}\to\mathbb{N}$ is called proper, if for every $k\in\mathbb{N}$, the set $\sigma^{-1}(\{k\})$ is finite. The set of all proper maps from $\mathbb{N}$ to itself is denoted by $\mathscr{S}$.
Remark 3.2.
A map $\sigma\colon\mathbb{N}\to\mathbb{N}$ is proper if and only if $\lim_{n\to\infty}\sigma(n)=\infty$, i.e., for every $N\in\mathbb{N}$, there exists $n_{0}\in\mathbb{N}$ such that $\sigma(n)>N$ for every $n\geq n_{0}$.
Now for a proper map $\sigma\colon\mathbb{N}\to\mathbb{N}$, we set
$S_{n}^{\sigma}=T_{\sigma(n)}\dots T_{\sigma(1)}.$
By $\lim_{n\to\infty}\sigma(n)=\infty$, $\lim_{n\to\infty}T_{\sigma(n)}=T$ (SOT) holds. Then we have the following analogue of [1, Proposition 2.3]. The proof is essentially the same, so we do not repeat it here.
Proposition 3.3.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$, and let $S_{n}^{\sigma}:=T_{\sigma(n)}\cdots T_{\sigma(1)}$. The following statements hold (WOT stands for the weak operator topology):
  • (1)
    $\lim_{n\to\infty}(S_{n}^{\sigma})^{*}=P$ (SOT). In particular, $\lim_{n\to\infty}S_{n}^{\sigma}=P$ (WOT) holds.
  • (2)
    Let $\xi\in H$. If the set $\{S_{n}^{\sigma}\xi\mid\,n\in\mathbb{N}\}$ is totally bounded, then $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi-P\xi\|=0$ holds.
  • (3)
    For every $\xi\in H$ and every $k\in\mathbb{N}$, $\lim_{n\to\infty}\|S_{n+k}^{\sigma}\xi-S_{n}^{\sigma}\xi\|=0$ holds.
Next, we introduce the $\sigma$-Paszkiewicz subspace $H_{\sigma}$.
Proposition 3.4.
Let $T_{1}\geq T_{2}\geq\dots$ be positive contractions on $H$ and let $\sigma\in\mathscr{S}$. Then the set
$H_{\sigma}:=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}\|S_{n}^{\sigma}\xi-P\xi\|=0\right\}=\left\{\xi\in H\,\middle|\,\lim_{n\to\infty}\|S_{n}^{\sigma}\xi\|=\|P\xi\|\right\}$
is a closed subspace of $H$ containing $P(H)$.
Proof.
It is clear that $H_{\sigma}$ is a vector subspace of $H$ and that $P(H)\subset H_{\sigma}$ because $S_{n}^{\sigma}P\xi=P\xi$ for every $\xi\in H$ and $n\in\mathbb{N}$. Let $\xi\in\overline{H_{\sigma}}$ and $\varepsilon>0$. Then there exists $\xi_{0}\in H_{\sigma}$ such that $\|\xi-\xi_{0}\|<\frac{\varepsilon}{3}$ holds. Since $\xi_{0}\in H_{\sigma}$, we may find an $n_{0}\in\mathbb{N}$ such that $\|S_{m}^{\sigma}\xi_{0}-S_{n}^{\sigma}\xi_{0}\|<\frac{\varepsilon}{3}$ for every $n,m\geq n_{0}$. Then for every $n,m\geq n_{0}$,
$\|S_{m}^{\sigma}\xi-S_{n}^{\sigma}\xi\|$ $\leq\|S_{m}^{\sigma}(\xi-\xi_{0})\|+\|S_{m}^{\sigma}\xi_{0}-S_{n}^{\sigma}\xi_{0}\|+\|S_{n}^{\sigma}\xi_{0}-S_{n}^{\sigma}\xi\|$
$<2\|\xi-\xi_{0}\|+\frac{\varepsilon}{3}<\varepsilon.$
Since $\varepsilon$ is arbitrary, this shows that $(S_{n}^{\sigma}\xi)_{n=1}^{\infty}$ is Cauchy, whence it converges to $P\xi$ by Proposition 3.3 (2). Therefore, $\xi\in H_{\sigma}$ holds. This shows that $H_{\sigma}$ is closed. Finally, to show the last equality, we show that for each $\xi\in H$ the following conditions are equivalent:
  • (a)
    $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi-P\xi\|=0.$
  • (b)
    $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi\|=\|P\xi\|.$
(a)$\implies$(b) is clear. Assume (b). Then because $\lim_{n\to\infty}S_{n}^{\sigma}=P$ (WOT), we have
$\|S_{n}^{\sigma}\xi-P\xi\|^{2}$ $=\|S_{n}^{\sigma}\xi\|^{2}+\|P\xi\|^{2}-2\,{\rm Re}\langle S_{n}^{\sigma}\xi,P\xi\rangle$
$\xrightarrow{n\to\infty}\|P\xi\|^{2}+\|P\xi\|^{2}-2\,{\rm Re}\langle P\xi,P\xi\rangle=0.$
Therefore, (a) holds. This finishes the proof. ∎
Then, inspired by the role played by the subspace “$Z$” in the work of Kopecká–Paszkiewicz [2, Lemma 3.4], we define the following closed subspace $H_{\mathscr{S}}$ of $H$:
Definition 3.5.
We define the $\sigma$-Paszkiewicz subspace $H_{\sigma}$ by
$H_{\sigma}=\{\xi\in H\mid\lim_{n\to\infty}S_{n}^{\sigma}\xi=P\xi\},$
and then we set
$H_{\mathscr{S}}=\bigcap_{\sigma\in\mathscr{S}}H_{\sigma}.$
Let $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$.
Definition 3.6.
We say that the generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if $H_{\mathscr{S}}=H$ holds.
Theorem 3.7.
The following statements hold.
  • (1)
    Both $H_{\mathscr{S}}$ is a closed subspace of $H$ which are invariant under all $T_{n}$.
  • (2)
    $e$ is a central projection in $\mathscr{M}$. In particular, if $\mathscr{M}$ is a factor, then the generalized Paszkiewicz conjecture holds for $T_{1}\geq T_{2}\geq\dots$ if and only if $H_{\mathscr{S}}\neq\{0\}$.
Proof.
(1) Let $\sigma\in\mathscr{S}$, $m,n\in\mathbb{N}$ and $\xi\in H_{\mathscr{S}}$. Then the map $\hat{\sigma}\colon\mathbb{N}\to\mathbb{N}$ defined by $\hat{\sigma}(1)=m,\hat{\sigma}(k+1)=\sigma(k),\,k\in\mathbb{N}$ is an element in $\mathscr{S}$. Thus,
$S_{n}^{\sigma}(T_{m}\xi)=S_{n+1}^{\hat{\sigma}}\xi\xrightarrow{n\to\infty}P\xi=P(T_{m}\xi)$
by $\xi\in H_{\hat{\sigma}}$. Therefore $T_{m}\xi\in H_{\mathscr{S}}$. This shows that $H_{\mathscr{S}}$ is invariant under $T_{m}$.
(2) By (1), $e$ belongs to $\mathscr{M}^{\prime}$. Assume that $f$ is a projection in $\mathscr{M}^{\prime}$ such that $e\sim f$ in $\mathscr{M}^{\prime}$ ($\sim$ denotes the Murray-von Neumann equivalence of projections). By $e\sim f$, there exists a partial isometry $u\in\mathscr{M}^{\prime}$ such that $u^{*}u=e$ and $uu^{*}=f$. Let $\xi\in f(H)$ and $\sigma\in\mathscr{S}$. Then $u^{*}\xi\in e(H)\subset H_{\sigma}$. By $P,S_{n}^{\sigma}\in\mathscr{M}$ and $u\in\mathscr{M}^{\prime}$, we have
$S_{n}^{\sigma}\xi=S_{n}^{\sigma}uu^{*}\xi=uS_{n}^{\sigma}(u^{*}\xi)\xrightarrow{n\to\infty}uPu^{*}\xi=Puu^{*}\xi=P\xi.$
This shows that $\xi\in H_{\sigma}$. Since $\sigma\in\mathscr{S}$ is arbitrary, we obtain $\xi\in H_{\mathscr{S}}=e(H)$. Therefore, $f\leq e$ holds. It then follows that
$e\leq\bigvee\left\{ueu^{*}\mid u\in U(\mathscr{M}^{\prime})\right\}\leq\bigvee\left\{f\in{\rm Proj}(\mathscr{M}^{\prime})\mid f\sim e\right\}\leq e,$
whence $e=\bigvee\left\{ueu^{*}\mid u\in U(\mathscr{M}^{\prime})\right\}=z(e)$ holds (here, $z(e)$ is the central support of $e$). The last claim is then immediate. ∎
We then give examples of sequences for which the generalized Paszkiewicz conjecture holds (that many of such examples exist was the motivation for this generalization). It is clear that the generalized Paszkiewicz conjecture is true when $\mathscr{M}=W^{*}(T_{1},T_{2},\dots)$ is a finite von Neumann algebra, because $\lim_{n\to\infty}(S_{n}^{\sigma})^{*}=P$ (SOT) for every $\sigma\in\mathscr{S}$ and the *-operation is SOT-continuous on the unit ball of $\mathscr{M}$. The next result shows that the generalized Paszkiewicz conjecture is true also for sequences with uniform spectral gap at 1, which generalizes [1].
Proposition 3.8.
If $T_{1}\geq T_{2}\geq\dots$ has uniform spectral gap at 1, then the generalized Paszkiewicz conjecture holds for it.
Lemma 3.9.
Let $T_{1}\geq T_{2}$ be positive contractions on $H$. Let $P_{i}=1_{\{1\}}(T_{i})\,(i=1,2)$. Then for each $i,j\in\{1,2\}$, $T_{i}P_{j}^{\perp}(H)\subset P_{\max\{i,j\}}^{\perp}(H)$ holds.
Proof.
The case $i=j$ is trivial. Assume $i<j$. Then $T_{i}\geq T_{j}$, whence $P_{i}\geq P_{j}$ and $T_{i}P_{j}=P_{j}T_{i}=P_{j}$. Thus $T_{i}P_{j}^{\perp}(H)=P_{j}^{\perp}T_{i}(H)\subset P_{j}^{\perp}(H)$. If $i>j$, then $P_{j}\geq P_{i}$, whence $P_{j}^{\perp}\leq P_{i}^{\perp}$. Therefore, $T_{i}P_{j}^{\perp}(H)\subset T_{i}P_{i}^{\perp}(H)\subset P_{i}^{\perp}(H)$. ∎
Proof of Proposition 3.8.
Let $\sigma\in\mathscr{S}$. Let $\varepsilon>0$ and $\xi\in P^{\perp}(H)$ be a unit vector. Then there exists $n_{0}\in\mathbb{N}$ such that $\|\xi-P_{n_{0}}^{\perp}\xi\|<\frac{1}{2}\varepsilon$ holds. By assumption, there exists $N\in\mathbb{N}$ and $\delta\in(0,1)$ such that $\sigma(T_{n})\cap(1-\delta,1)=\emptyset$ for every $n\geq N$. We may assume that $N\geq n_{0}$. Using the properness of $\sigma$, we may find an increasing sequence of natural numbers
$2\leq n_{1}<n_{2}<\cdots$
such that $\sigma(n_{k})>\max\{N,\sigma(1),\dots,\sigma(n_{k}-1)\}$ for every $k\in\mathbb{N}$. Indeed, by the properness of $\sigma$, there exists $n_{1}\geq 2$ such that $\sigma(n_{1})\geq N$ holds. Assume that we have found $n_{1}<n_{2}<\cdots<n_{k}$. Then by the properness of $\sigma$, there exists $n\in\mathbb{N}$ for which $n>n_{k}$ and $\sigma(n)>\sigma(n_{k})$ holds. Let $n_{k+1}$ be the smallest such $n$. Then $\sigma(n_{k+1})>N,\sigma(n_{k})$, and if $n<n_{k}$, then $\sigma(n)<\sigma(n_{k})<\sigma(n_{k+1})$ and if $n_{k}\leq n<n_{k+1}$, then $\sigma(n_{k+1})>\sigma(n_{k})\geq\sigma(n)$ by the choice of $n_{k+1}$. By induction, we have the $n_{1}<n_{2}<\dots$ with the required properties. Choose $k\in\mathbb{N}$ for which $(1-\delta)^{k}<\frac{1}{2}\varepsilon$ holds. Then for every $n\geq n_{k}$, we have
$\|S_{n}^{\sigma}\xi\|\leq\|S_{n_{k}}^{\sigma}\xi\|\leq\|S_{n_{k}}^{\sigma}(\xi-P_{n_{0}}^{\perp}\xi)\|+\|S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
and by a repeated use of Lemma 3.9, we have
$S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi=T_{n_{k}}(\underbrace{T_{\sigma(n_{k}-1)}\cdots T_{\sigma(1)}P_{n_{0}}^{\perp}\xi)}_{\in P_{\max\{\sigma(1),\dots,\sigma(n_{k}-1),n_{0}\}}^{\perp}(H)}$
And by $\sigma(n_{k})>\max(\sigma(1),\dots,\sigma(n_{k}-1),n_{0},N)$, we have
$P_{\max\{\sigma(1),\dots,\sigma(n_{k}-1),n_{0}\}}^{\perp}(H)\subset P_{\sigma(n_{k})}^{\perp}(H)=1_{[0,1-\delta]}(T_{\sigma(n_{k})})(H).$
This implies that
$\|S_{n_{k}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$ $\leq(1-\delta)\|T_{\sigma(n_{k}-1)}\cdots T_{\sigma(n_{k-1}+1)}S_{n_{k-1}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq(1-\delta)\|S_{n_{k-1}}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq\dots\leq(1-\delta)^{k}\|S_{n_{1}-1}^{\sigma}P_{n_{0}}^{\perp}\xi\|$
$\leq(1-\delta)^{k}<\frac{\varepsilon}{2}.$
Therefore $\|S_{n}^{\sigma}\xi\|<\varepsilon\,(n\geq n_{k})$. Since $\varepsilon$ is arbitrary, we obtain $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi\|=0$ for every $\xi\in P^{\perp}(H)$, i.e., $H_{\sigma}=H$ holds. ∎
Proposition 3.10.
Let $T_{1}\geq T_{2}\geq\cdots$ be a decreasing sequence of positive contractions on $H$. Assume that $1$ is isolated in $\sigma(T)$ and $\lim_{n\to\infty}\|T_{n}-T\|=0$. Then $\lim_{n\to\infty}\|S_{n}^{\sigma}-P\|=0$ for every $\sigma\in\mathscr{S}$. In particular, the generalized Paszkiewicz conjecture is true for $T_{1}\geq T_{2}\geq\cdots$.
Proof.
By assumption, there exists $\delta\in(0,1)$ such that $\sigma(T)\cap(1-\delta,1)=\emptyset$. In particular, $\|T^{k}-P\|=\|T^{k}P^{\perp}\|\leq(1-\delta)^{k}\xrightarrow{k\to\infty}0$. Therefore, for a given $\varepsilon>0$, there exists $k\in\mathbb{N}$ such that $\|T^{k}-P\|<\varepsilon$. Then by $T_{j}P=P\,(j\in\mathbb{N})$ and $\lim_{n\to\infty}\|T_{\sigma(n+1)}\dots T_{\sigma(n+k)}-T^{k}\|=0$, we have
$\limsup_{n\to\infty}\|(S_{n}^{\sigma})^{*}-P\|$ $=\limsup_{n\to\infty}\|(S_{n+k}^{\sigma})^{*}-P\|$
$=\limsup_{n\to\infty}\|T_{\sigma(1)}\dots T_{\sigma(n)}(T_{\sigma(n+1)}\dots T_{\sigma(n+k)}-P)\|$
$\leq\limsup_{n\to\infty}\|T_{\sigma(n+1)}\dots T_{\sigma(n+k)}-P\|$
$=\|T^{k}-P\|<\varepsilon.$
Since $\varepsilon$ is arbitrary, we obtain $\lim_{n\to\infty}\|(S_{n}^{\sigma})^{*}-P\|=\lim_{n\to\infty}\|S_{n}^{\sigma}-P\|=0$. ∎
Remark 3.11.
If $T_{1}\geq T_{2}\geq\dots$ is a decreasing sequence of positive contraction as in Proposition 3.10, then it has uniform spectral gap at 1 (this was pointed out by Hiroki Matui). Thus, Proposition 3.10 is a special case of Proposition 3.8. On the other hand, note that for a general decreasing sequence of positive contractions with uniform spectral gap at 1, one cannot hope that $S_{n}$ converges in norm (consider e.g., the case where all $T_{n}$ are projections).
If $T_{n}$ converges to $T$ faster (but $1$ is possibly not isolated in $\sigma(T)$), then we also have a slightly weaker conclusion.
Proposition 3.12.
Let $T_{1}\geq T_{2}\geq\cdots$ be a sequence of positive contractions on $H$. Assume that $\sum_{n=1}^{\infty}\|T-T_{n}\|<\infty$ holds. Then for each $\sigma\in\mathscr{S}$ such that $\sup_{n\in\mathbb{N}}\sharp\sigma^{-1}(\{n\})<\infty$, $\lim_{n\to\infty}S_{n}^{\sigma}=P$ (SOT) holds.
Proof.
Let $\xi\in H$ be a unit vector and $\varepsilon>0$. By the hypothesis on $\sigma$, we have $\sum_{n=1}^{\infty}\|T-T_{\sigma(n)}\|<\infty$. Choose $n_{0}\in\mathbb{N}$ such that $\sum_{n=n_{0}+1}^{\infty}\|T-T_{\sigma(n)}\|<\frac{\varepsilon}{2}$ holds. Since $\lim_{n\to\infty}T^{n}=P$ (SOT), there exists $n_{1}\in\mathbb{N}$ such that $\|T^{n}S_{n_{0}}^{\sigma}\xi-PS_{n_{0}}^{\sigma}\xi\|=\|T^{n}S_{n_{0}}^{\sigma}\xi-P\xi\|<\frac{\varepsilon}{2}$ for every $n\geq n_{1}$.
Let $S_{n,n_{0}+1}^{\sigma}=T_{\sigma(n+n_{0})}\cdots T_{\sigma(n_{0}+1)}$ ($n$ products). For every $n\in\mathbb{N}$, we have
$\|S_{n,n_{0}+1}^{\sigma}-T^{n}\|$ $\leq\sum_{k=1}^{n-1}\|T_{\sigma(n+n_{0})}\cdots T_{\sigma(n_{0}+k+1)}(T_{\sigma(n_{0}+k)}-T)T^{k-1}\|+\|(T_{\sigma(n+n_{0})}-T)T^{n-1}\|$
$\leq\sum_{k=n_{0}+1}^{\infty}\|T-T_{\sigma(k)}\|<\frac{\varepsilon}{2}.$
Therefore, by $S_{n+n_{0}}^{\sigma}=S_{n,n_{0}+1}^{\sigma}S_{n_{0}}^{\sigma}$, it holds that for every $n\geq n_{1}$,
$\|S_{n+n_{0}}^{\sigma}\xi-P\xi\|$ $\leq\|(S_{n,n_{0}+1}^{\sigma}-T^{n})S_{n_{0}}^{\sigma}\xi\|+\|T^{n}S_{n_{0}}^{\sigma}\xi-PS_{n_{0}}^{\sigma}\xi\|$
$<\|S_{n,n_{0}+1}^{\sigma}-T^{n}\|+\frac{\varepsilon}{2}<\varepsilon.$
Since $\varepsilon$ is arbitrary, this shows that $\lim_{n\to\infty}\|S_{n}^{\sigma}\xi-P\xi\|=0$. ∎
Acknowledgments
We would like to thank Professors Masaru Nagisa and Mitsuru Uchiyama for useful discussions regarding the spectral order which we used in an earlier version of the paper and for pointing us to related literatures, Professor Hiroki Matui for his useful comment at Remark 3.11. H. Ando is supported by Japan Society for the Promotion of Sciences (JSPS) KAKENHI 20K03647. N. Ozawa is partially supported by JSPS KAKENHI 20H00114 and 24K00527.